### v3

# DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection

Abstract
The rapid development of multilingual large language models (LLMs) highlights the need for high-quality, diverse, and clean multilingual datasets. In this paper, we introduce DCAD-2000 (Data Cleaning as Anomaly Detection), a large-scale multilingual corpus built using newly extracted Common Crawl data and existing multilingual datasets. DCAD-2000 includes over 2,282 languages, 46.72TB of data, and 8.63 billion documents, spanning 155 high- and medium-resource languages and 159 writing scripts. To overcome the limitations of current data cleaning methods, which rely on manual heuristic thresholds, we propose reframing data cleaning as an anomaly detection task. This dynamic filtering approach significantly enhances data quality by identifying and removing noisy or anomalous content. By fine-tuning LLMs on DCAD-2000, we demonstrate the effectiveness of the data quality improvements, the robustness of the data cleaning pipeline, and enhanced performance for low-resource languages across multiple multilingual benchmarks.
[Uncaptioned image] Dataset: https://huggingface.co/datasets/openbmb/DCAD-2000
[Uncaptioned image] Pipeline: https://github.com/yl-shen/DCAD-2000
1 Introduction
Large language models (LLMs) have achieved great progress on a variety of NLP tasks by leveraging vast amounts of training data [1]. However, their performance remains heavily biased towards high-resource languages [2, 3]. To improve the multilingual capabilities of LLMs, a common strategy is to incorporate large amounts of non-English data, either by continue pretraining [4] or by instruction tuning in multilingual settings [5]. Therefore, constructing large-scale, high-quality multilingual datasets is crucial for enhancing the multilingual performance of LLMs.
Table 1: Comparison of multilingual datasets constructed from Common Crawl ($C$) and our constructed DCAD-2000, focusing on the latest $C$ version used, the total number of languages supported, distribution across resource categories (high, medium, low, very low), and training readiness. The $C$ version marked with underline indicates an inferred version due to the lack of explicit specification in the original paper. The “Training-Ready” column indicates whether the dataset is ready for training LLMs without requiring further data cleaning.
Dataset CC Version
#Langs
#Langs
#Langs
#Langs
#Langs
(very low)
Training-Ready
mC4 [6] $C$-MAIN-2020-34 101 0 43 52 6
OSCAR 23.01 [7] $C$-MAIN-2022-49 153 6 42 25 80
Glot500 [8] $C$-MAIN-2020-34 511 0 108 79 324
CulturaX [9] $C$-MAIN-2022-49 167 11 47 27 82
Madlad-400 [10] $C$-MAIN-2022-33 419 7 46 39 327
MaLA [11] $C$-MAIN-2022-49 939 1 125 78 735
Glotcc [12] $C$-MAIN-2023-50 1331 0 10 52 1269
HPLT-v1.2 [13] $C$-MAIN-2022-40 191 12 53 38 88
Fineweb-2 [14] $C$-MAIN-2024-18 1915 10 62 49 1794
DCAD-2000 $C$-MAIN-2024-46 2282 13 142 124 2003
Recent efforts to construct multilingual datasets have built several notable resources, including CulturaX [9], HPLT [13], Madlad-400 [10], MaLA [15], and Glotcc [12], which cover 167, 191, 419, 939, and 1,331 languages, respectively. While these datasets have made significant contributions, they exhibit three major limitations, as summarized in Table 1: (1) Outdated data sources: These datasets primarily rely on older versions of Common Crawl [^1] https://commoncrawl.org, which results in outdated knowledge and an elevated risk of hallucination [16]. (2) Limited coverage of high- and medium-resource languages [^2] We follow the criteria from Flores-101 [17] to categorize languages: High: >100M>100M; Medium: $(1M,100M)$; Low: $(100K,1M)$; Very Low: <100K<100K.: For instance, Fineweb-2 [14], which supports 1,915 languages, includes data from only 10 high-resource languages and 62 medium-resource languages. (3) Insufficient data cleaning: Despite being cleaned, recent studies [18, 19] indicate that these datasets still contain a significant amount of noise, which makes them difficult to directly employ in training multilingual LLMs. For instance, Sailor [18] founds that 31.11% of the data in Madlad-400 could still be further removed by advanced data cleaning.
Traditional data cleaning workflows [20] often rely on document-level features (e.g., language identification; 21) and fixed thresholds to filter out low-quality data. However, this approach struggles with cross-lingual consistency due to feature distribution differences. For example [^1] Please refer to Appendix A for more details., the variation in average word counts across languages, datasets, and shards demonstrates the limitations of heuristic threshold-based data cleaning methods. Notably, while Fineweb-2 fine-tunes thresholds for more than 1,000 languages, this process is computationally intensive and time-consuming.
To address these challenges, we introduce DCAD-2000, a new large-scale, high-quality multilingual dataset that can be directly applied to LLM training. DCAD-2000 covers 2282 languages (155 high/medium languages), incorporating the latest Common Crawl data (November 2024; CC-MAIN-2024-46) and existing multilingual datasets. Additionally, we propose a novel language-agnostic data cleaning approach that treats data cleaning as an anomaly detection [22] problem, distinguishing it from traditional threshold-based methods [14, 23]. Our approach extracts eight statistical features from each document to evaluate quality, including number of words, character/word repetition ratio, special character/word ratio, stopword ratio, flagged words ratio, language identification score and perplexity score. Anomaly detection algorithms dynamically identify and remove outliers by recognizing deviations from typical document quality metrics.
We conduct a comprehensive analysis of DCAD-2000 with respect to document distribution, linguistic and geographical characteristics, writing scripts, and resource classification (Section 4). By fine-tuning LLMs on DCAD-2000, we validate the effectiveness of its data quality and data cleaning pipeline. Furthermore, we demonstrate the superiority of DCAD-2000 across various language categories (high, medium, low and very low) in multiple multilingual benchmarks, including SIB-200 [24], Glot500 [8] and FLORES-200 [25] (Section 5).
In summary, we make the following contributions: (1) We propose a novel data cleaning framework that frames the task as anomaly detection, offering a language-agnostic and adaptive solution without manual threshold tuning. (2) We release DCAD-2000, a comprehensive multilingual dataset covering over 2,282 languages, containing 8.63B of documents, 46.72TB of disk size and 159 writing scripts with metadata annotations. (3) Extensive evaluation across multiple multilingual benchmarks demonstrates the effectiveness of both the data quality and the data cleaning pipeline.
2 Related Work
Multilingual Dataset for Pretraining.
Enhancing the multilingual capabilities of LLMs often involves continuing pretraining on large-scale multilingual datasets [11, 15]. These datasets can be broadly categorized into curated corpora, domain-specific corpora, and web-crawled corpora. (I) Curated Corpora. Curated datasets are carefully gathered by experts from high-quality sources such as books [23], academic publications [26], and encyclopedia entries [27, 28, 29]. (II) Domain-Specific Corpora. In addition to general-domain data, fine-tuning LLMs on domain-specific datasets is crucial for improving performance in specialized domains like finance [30], healthcare [31], legal [32], and education [33, 34]. (III) Web-Crawled Corpora. Web-crawled datasets, particularly those derived from Common Crawl, provide large-scale multilingual coverage by leveraging an open repository of over 250 billion web pages. These datasets include mC4 [6], CC-100 [35], OSCAR [7], Glotcc [12], Fineweb [36], and Fineweb-2 [14]. While curated and domain-specific corpora offer high-quality content with limited language coverage, web-crawled corpora provide broader multilingual coverage but often suffer from noise and lower data quality [18, 19].
Data Cleaning.
Data cleaning is an essential step in preparing high-quality datasets for training robust LLMs. It involves filtering noisy, irrelevant, or harmful content and can be broadly classified into model-based and heuristic-based approaches [37]. (I) Model-Based Methods. Model-based approaches employ classifiers or LLMs to distinguish between high-quality and low-quality data. For instance, content safety models [38] filter out explicit or gambling-related content, while quality classifiers remove low-relevance text [39]. LLM-based methods focus on generating prompts for cleaning [40] or integrating error detection and correction into the pipeline [41, 42]. (II) Heuristic-Based Methods. Heuristic approaches apply predefined rules to filter content at both document and sentence levels. At the document level, strategies include filtering by language identification scores or scoring documents with language models [9, 23]. At the sentence level, rules are applied to remove incomplete or irrelevant content, such as HTML tags or excessively short sentences [6, 7]. While model-based methods offer high precision but face scalability challenges, heuristic-based methods are more efficient yet less adaptable to diverse multilingual data.
3 DCAD-2000
To overcome the limitations of existing multilingual datasets, we introduce DCAD-2000, a large-scale, high-quality multilingual dataset constructed by integrating data from latest version of Common Crawl and existing multilingual datasets (Section 3.1). This dataset is cleaned using our proposed novel framework, which treats data cleaning as an anomaly detection problem (Section 3.2). The construction of DCAD-2000 is supported by robust computational resources, as detailed in Section 3.3.
3.1 Data Collection
To ensure comprehensive and robust multilingual data representation, DCAD-2000 integrates data from four main sources: MaLA, Fineweb, Fineweb-2, and newly extracted Common Crawl data. Each source is selected based on its unique contribution to multilingual coverage, data quality, and freshness, with careful consideration to complementarity to minimize redundancy. Specifically, MaLA and Fineweb-2 are prioritized due to their broad language coverage and high-quality curation, which complements other widely used datasets like mC4 [6] and OSCAR [7].
MaLA Corpus [11]. The MaLA corpus covers 939 languages, aggregating data from diverse sources including Bloom [43], CC100 [35], Glot500 [8], among others. Deduplication is performed using MinHashLSH [44], which is particularly effective in removing near-duplicate entries that often arise from common web sources. Language codes are based on ISO 639-3222https://en.wikipedia.org/wiki/ISO_639-3 standards, and language-specific scripts are supported by GlotScript [^3] https://github.com/cisnlp/GlotScript.
Fineweb Corpus [36]. Fineweb is a high-quality English web dataset extracted from Common Crawl, consisting of over 15 trillion tokens and updated monthly. Data cleaning and deduplication are performed using the Datatrove library. [^4] https://github.com/huggingface/datatrove For DCAD-2000, we incorporate data from the November 2024 release (CC-MAIN-2024-46) to ensure freshness and up-to-date relevance of the data.
Fineweb-2 Corpus [14]. Fineweb-2 expands Fineweb to include multilingual data, covering 1,915 languages. It processes 96 Common Crawl dumps from 2013 (CC-MAIN-2013-20) to April 2024 (CC-MAIN-2024-20). The deduplication process within Fineweb-2 is similarly handled using the Datatrove library, ensuring the exclusion of redundant entries and maintaining high-quality multilingual coverage.
Newly Extracted Common Crawl Data. To incorporate the most recent multilingual data, we extract and process Common Crawl dumps from May 2024 (CC-MAIN-2024-22) to November 2024 (CC-MAIN-2024-46). Using the Fineweb-2 pipeline [^5] https://github.com/huggingface/fineweb-2, we process 21.54TB of multilingual data, ensuring that the data remains fresh and suitable for downstream tasks. This further extends the multilingual data pool and enhances the coverage across underrepresented languages.
3.2 Data Cleaning as Anomaly Detection
Traditional data cleaning methods rely on fixed thresholds for document-level features, making them less adaptable to the diversity of multilingual data. To address this, we propose a novel framework that formulates data cleaning as an anomaly detection task, which involves feature extraction (Section 3.2.1) and anomaly detection (Section 3.2.2).
3.2.1 Feature Extraction
Inspired by Roots [23] and CulturaX [9], we extract eight statistical features from each document to evaluate text quality. Each feature is selected for its ability to capture important characteristics of the text, contributing to robust anomaly detection. Let $t$ represent a document; the extracted features are: (1) Number of Words, $n_{w}(t)$: Total token count after tokenization. (2) Character Repetition Ratio, $r_{c}(t)$: Fraction of repeated character sequences, highlighting noise or encoding errors. (3) Word Repetition Ratio, $r_{w}(t)$: Proportion of repeated words, indicating redundancy or unnatural text. (4) Special Characters Ratio, $r_{s}(t)$: Fraction of special characters based on language-specific lists from ROOTs Corpus [23]. These lists include emojis, whitespace types, numbers, and punctuation. (5) Stopwords Ratio, $r_{\text{stop}}(t)$: Proportion of stopwords using Fineweb-2’s language-specific stopword lists. (6) Flagged Words Ratio, $r_{\text{flag}}(t)$: Fraction of toxic or profane words from Toxicity-200 [25] or public repositories [^6] https://github.com/thisandagain/washyourmouthoutwithsoap. (7) Language Identification (LID) Score, $s_{\text{lid}}(t)$: Confidence score from GlotLID [21], supporting over 2,000 languages. (8) Perplexity Score, $s_{\text{ppl}}(t)$: Using KenLM [45], we train a language model for each language on multilingual Wikipedia data (November 2023) [^7] For languages not covered by the trained models, a default perplexity score of 500 is assigned..
The feature vector for each document is defined as:
$\mathbf{x}=\left[n_{w}(t),\,r_{c}(t),\,r_{w}(t),\,r_{s}(t),\,r_{\text{stop}}(t),\,r_{\text{flag}}(t),\,s_{\text{lid}}(t),\,s_{\text{ppl}}(t)\right]^{\top}\in\mathbb{R}^{8}.$ (1)
3.2.2 Anomaly Detection
After extracting feature vectors $\mathbf{x}\in\mathbb{R}^{8}$, we standardize each feature to handle differences in scale. The standardized value $\tilde{x}_{j}$ for the $j$-th feature is given by:
$\tilde{x}_{j}=\frac{x_{j}-\mu_{j}}{\sigma_{j}},\quad j=1,\ldots,8.$ (2)
where $\mu_{j}$ and $\sigma_{j}$ are the mean and standard deviation of the $j$-th feature across the dataset. The standardized feature vector is:
$\tilde{\mathbf{x}}=\frac{\mathbf{x}-\boldsymbol{\mu}}{\boldsymbol{\sigma}}.$ (3)
where $\boldsymbol{\mu}=[\mu_{1},\mu_{2},\ldots,\mu_{8}]^{\top}$ and $\boldsymbol{\sigma}=[\sigma_{1},\sigma_{2},\ldots,\sigma_{8}]^{\top}$ are the vectors of means and standard deviations, respectively.
[…]
$\phi(\tilde{\mathbf{x}})=2^{-\frac{h(\tilde{\mathbf{x}})}{c(n)}}.$ (4)
[…]
$c(n)=2H(n-1)-\frac{2(n-1)}{n}.$ (5)
where $H(i)$ is the $i$-th harmonic number, defined as $H(i)=\sum_{k=1}^{i}\frac{1}{k}$.
An anomaly score $\phi(\tilde{\mathbf{x}}):\mathbb{R}^{8}\to\mathbb{R}$ is defined to quantify how far a document deviates from typical data. Higher scores indicate a higher likelihood of anomalies. To classify a document, we use the decision rule:
[…]
where $\tau\in\mathbb{R}$ is a hyperparameter determined empirically or through cross-validation. [^9] We use the default setting of the specific anomaly detection algorithm.
[…]
$\mathcal{X}_{\text{keep}}$ $=\{\tilde{\mathbf{x}}\in\tilde{\mathcal{X}}:f(\tilde{\mathbf{x}})=1\},$ (7)
$\mathcal{X}_{\text{remove}}$ $=\{\tilde{\mathbf{x}}\in\tilde{\mathcal{X}}:f(\tilde{\mathbf{x}})=-1\}.$ (8)
Following anomaly detection, the dataset is partitioned into a clean subset $\mathcal{X}_{\text{keep}}$ and an anomalous subset $\mathcal{X}_{\text{remove}}$. The former is retained for downstream tasks such as model training, while the latter may be discarded or further examined for potential data quality issues.
Figure 1: Scatter plots of eight features extracted from a Chinese corpus during the data cleaning process, with data points color-coded according to their anomaly labels. The yellow points represent high-quality data, while the purple points indicate low-quality data.
3.2.3 Visualization
To qualitatively evaluate the separation achieved by our data cleaning framework, we present scatter plots of the eight feature dimensions in Figure 1, with data points color-coded by their anomaly labels. These visualizations facilitate the interpretation of decision boundaries and highlight the features that contribute most significantly to the detection process. We observe well-defined clusters separating anomalous and non-anomalous data points, with anomalies exhibiting distinct patterns compared to the majority of the data. Features such as the language identification score ($s_{\text{lid}}(t)$) and perplexity score ($s_{\text{ppl}}(t)$) are expected to be particularly discriminative in identifying anomalies, as they capture linguistic irregularities and unexpected text patterns. For example, low lidlid or unusually high pplppl scores often indicate problematic text, such as spam, low-quality content, or noise. The framework effectively identifies and removes such low-quality text samples, which can be easily visualized by the separation of these points in the scatter plots.
3.3 Computational Resources
The construction of the DCAD-2000 dataset leveraged Ksyun servers [^10] https://k8s.ksyun.com to process and clean the multilingual data efficiently. Each server instance is equipped with 32 CPU cores, 128GB of memory, and 100GB of disk storage, which is utilized for intermediate data handling and memory-intensive operations such as anomaly detection. The workload is managed using container orchestration tools, Kubernetes [^11] https://kubernetes.io, with up to 100 parallel tasks running per job to ensure scalability.
4 Dataset Analysis
(a) Document Distribution
(b) Geographical Distribution
(c) Script Distribution
Figure 2: Document distribution and linguistic diversity in DCAD-2000.
In this section, we analyze the characteristics of DCAD-2000, focusing on document distribution across sources, geographic and script coverage, resource categorization of languages, and the effect of data cleaning on dataset size and quality.
Document Distribution Across Data Sources. The DCAD-2000 dataset is derived from four primary sources: MaLA, Fineweb, Fineweb-2, and Newly Extracted Common Crawl data (New CC), as described in Section 3.1. Figure 2(a) presents the distribution of documents across these sources, with Fineweb-2 and New CC collectively contributing 47.5% and 39.3% of the total dataset, respectively. These two sources play a significant role in ensuring the dataset’s emphasis on both language diversity (Fineweb-2) and corpus freshness (New CC). MaLA, though contributing 11.1% of the total dataset, brings in valuable content from non-Common Crawl sources, further enriching the diversity of the dataset, especially for low-resource languages.
Geographical Coverage of Languages. Figure 2(b) shows the geographical distribution of languages in DCAD-2000, based on the number of unique languages available in each region, as classified by Glottolog. [^12] Geographic data source: https://glottolog.org The dataset spans languages from all major world regions, with the largest proportions originating from Africa (28.6%), Papunesia (26.3%) and Eurasia (23.8%). This coverage ensures robust support for multilingual applications across varied regional contexts, including densely populated areas like Eurasia and sparsely populated regions such as Papunesia and Australia. While Eurasia is more heavily represented, this diversity of linguistic coverage helps ensure that the dataset remains useful for training LLMs in diverse regional environments.
Script Distribution. Figure 2(c) illustrates the distribution of languages in DCAD-2000 by writing system. The dataset supports 159 scripts, with the Latin script dominating at 79.4%, followed by Cyrillic (3.9%), Arabic (2.6%), and Devanagari (2.1%), among others. This diversity in scripts enables a wide range of cross-lingual and script-specific tasks. However, the inclusion of minority scripts, especially those with limited resources, poses unique challenges, such as optical character recognition (OCR) difficulties for certain scripts or inconsistent text quality. Despite these challenges, DCAD-2000 ensures comprehensive coverage by including data from diverse scripts. A complete list of supported scripts is provided in Appendix B.
Language Resource Classification. Following the classification approach proposed by Flores [17], we categorize languages in DCAD-2000 into four groups based on corpus size: high-resource, medium-resource, low-resource, and extremely low-resource. Table 1 shows the distribution across these categories. The dataset includes 155 high- and medium-resource languages, while low-resource languages make up a significant portion, which reflects DCAD-2000’s commitment to supporting underrepresented languages. Notably, DCAD-2000 surpasses other corpora in its balance between high-resource and low-resource languages, which can have a significant impact on multilingual model training. The distribution of languages across categories ensures that the dataset is well-suited for developing models that perform effectively across diverse language resources.
Impact of Data Cleaning. We summarize the document count, token count, and disk size of the high/medium/low resource languages in DCAD-2000 before and after the data cleaning process. Complete details are provided in Appendix C. The cleaning process results in the removal of a substantial amount of noisy data, even from datasets like MaLA, Fineweb, and Fineweb-2, which had already been subject some cleaning. This aligns with the findings from [18, 19]. For example, in the MaLA dataset, 8.05 million documents are removed for the hbs_Latn language, which suggests the necessity of rigorous data cleaning to enhance dataset quality. Overall, the cleaning process removed approximately 7.69% of the documents across all languages, significantly improving the quality of the dataset by reducing noise and increasing relevance for model training (Section 5).
5 Evaluation
Following Fineweb-2 [14], we conduct a series of experiments on the FineTask benchmark [^13] https://huggingface.co/spaces/HuggingFaceFW/blogpost-fine-tasks to evaluate the effectiveness of our proposed data cleaning pipeline and assess the quality of the DCAD-2000 dataset. FineTask comprises tasks in nine languages—Chinese, French, Arabic, Russian, Thai, Hindi, Turkish, Swahili, and Telugu—and covers a diverse set of NLP tasks, including reading comprehension, commonsense reasoning, natural language understanding, and text generation. To investigate the impact of different data cleaning strategies and anomaly detection algorithms, we continue pretraining on three typical LLMs: LLaMA-3.2-1B[47], Qwen-2.5-7B[48], and Aya-expanse-32B[49]. Additionally, we analyze the performance across different resource categories using the SIB-200[24], Glot500-c [8], and FLORES-200 [25] benchmarks. We report normalized accuracy for FineTask, raw accuracy for SIB-200, negative log-likelihood (NLL) for Glot500-c, and BLEU scores for FLORES-200. Full experimental settings and results are provided in Appendix D.
Figure 3: The performance comparison of models trained using various data cleaning methods.
Impact of Different Data Cleaning Strategies. We evaluate the effectiveness of our proposed anomaly detection-based data cleaning framework by comparing model performance across various cleaning strategies. As illustrated in Figure 3, the baseline model trained on raw, unfiltered data consistently underperforms relative to all cleaning methods. This performance gap is primarily due to noisy, irrelevant, or inconsistent data that hinders model generalization. Traditional threshold-based filtering [^14] We use the implementation from https://github.com/bigscience-workshop/data-preparation, which removes low-quality samples using fixed rules based on features, yields modest improvements. In contrast, our anomaly detection-based approach dynamically identifies and filters anomalous or noisy data, resulting in significantly enhanced model performance. Models trained using our method achieve normalized accuracy improvements of approximately 5–20% over the baseline, and outperform the threshold-based approach by 3–10%. While the threshold-based method is more computationally efficient due to its simplicity, our framework—though more resource-intensive—effectively captures subtle data issues that static rules often overlook.
Table 2: The performance of various anomaly detection algorithms. Bold and underlined numbers indicates the best and second-best results respectively.
LLaMA-3.2-1B Qwen-2.5-7B Aya-expanse-32B
Baseline Iso_Forest OC_SVM LOF K-Means Baseline Iso_Forest OC_SVM LOF K-Means Baseline Iso_Forest OC_SVM LOF K-Means
Arabic 0.07 0.21 0.18 0.21 0.14 0.63 0.71 0.68 0.65 0.68 0.69 0.75 0.70 0.71 0.69
Turkish 0.07 0.27 0.29 0.17 0.15 0.65 0.72 0.73 0.67 0.68 0.75 0.79 0.77 0.76 0.77
Swahili 0.08 0.29 0.25 0.19 0.19 0.25 0.34 0.27 0.35 0.27 0.35 0.44 0.36 0.37 0.41
Russian 0.10 0.24 0.19 0.18 0.15 0.74 0.79 0.75 0.75 0.76 0.77 0.82 0.79 0.80 0.79
Telugu 0.02 0.06 0.05 0.04 0.04 0.16 0.24 0.26 0.20 0.21 0.15 0.25 0.19 0.21 0.27
Thai 0.14 0.21 0.18 0.18 0.15 0.57 0.64 0.59 0.59 0.61 0.38 0.46 0.42 0.43 0.40
Chinese 0.12 0.32 0.28 0.25 0.21 0.75 0.82 0.77 0.76 0.78 0.69 0.75 0.71 0.71 0.73
French 0.11 0.35 0.37 0.30 0.23 0.74 0.80 0.76 0.76 0.75 0.74 0.79 0.76 0.76 0.76
Hindi 0.07 0.21 0.17 0.16 0.14 0.49 0.57 0.52 0.53 0.52 0.69 0.74 0.72 0.73 0.72
Comparison of Anomaly Detection Algorithms. We compare several classical anomaly detection algorithms to identify the most effective approach for constructing DCAD-2000. The evaluated methods include Isolation Forest (ISO_Forest;46), One-Class SVM (OC_SVM; 50), Local Outlier Factor (LOF;51), and K-Means [52], using implementations from scikit-learn [^15] https://scikit-learn.org. We provide the comparison of different algorithms in Appendix D.3. Table 2 reports the performance of these algorithms in cleaning the dataset. While all anomaly detection methods outperform the unfiltered baseline, the performance of OC_SVM, LOF, and K-Means is notably inconsistent. These algorithms often require extensive parameter tuning—such as selecting the number of neighbors for LOF or the kernel type for OC_SVM, which introduces sensitivity to hyperparameters and increases computational overhead. In contrast, ISO_Forest demonstrates more stable and robust performance across experiments, attributed to its efficiency in handling noisy, high-dimensional multilingual data. Unlike other methods, ISO_Forest delivers reliable results without intensive hyperparameter tuning, making it particularly suitable for large-scale multilingual datasets. However, ISO_Forest can be more computationally demanding than simpler methods like K-Means, especially in high-dimensional settings (our feature vectors have eight dimensions, as described in Section 3.2.2). Despite this trade-off, its robustness and scalability establish ISO_Forest as the most appropriate choice for data cleaning in DCAD-2000.
(a) French – LLaMA-3.2-1B
(b) Chinese – Qwen-2.5-7B
(c) Turkish – Aya-expanse-32B
Figure 4: Comparison of DCAD-2000 with existing multilingual corpora for three languages—French, Chinese, and Turkish—evaluated using different multilingual LLMs.
Comparison with Other Multilingual Datasets. To validate the quality of DCAD-2000, we compare it against existing multilingual corpora on the FineTask benchmark. These corpora include datasets constructed from New CC, MaLA, and Fineweb-2 as described in Section 3.1. As shown in Figure 4, models trained on DCAD-2000 consistently outperform those trained on other datasets, achieving higher normalized accuracy. The improvements can be attributed to the enhanced data quality, diversity, and reduced noise resulting from our comprehensive cleaning pipeline. Specifically, DCAD-2000 provides greater linguistic diversity and a more balanced representation of low-resource languages, leading to improved performance on tasks involving underrepresented languages like Swahili and Telugu.
Table 3: Performance across different language categories. We use accuracy ($\uparrow$) in SIB-200, negative log-likelihood ($\downarrow$) in Glot500-c and BLEU ($\uparrow$) in FLORES-200. Improvements are highlighted accordingly.
LLaMA-3.2-1B Qwen-2.5-7B Aya-expanse-32B
Fineweb-2 New CC DCAD-200 Fineweb-2 New CC DCAD-200 Fineweb-2 New CC DCAD-200
SIB-200 ($\uparrow$) H 8.24 8.86 10.37 ↑2.13 33.41 34.53 38.26 ↑4.85 41.72 42.41 47.93 ↑6.21
M 7.31 7.92 9.15 ↑1.84 28.72 29.86 32.65 ↑3.93 32.25 33.39 38.16 ↑5.91
L 6.06 6.45 7.83 ↑1.77 23.58 24.22 27.12 ↑3.54 26.87 27.57 33.24 ↑6.37
VL 3.68 4.27 5.24 ↑1.56 13.25 15.43 21.57 ↑8.32 17.23 19.5 26.38 ↑9.15
Glot500-c test ($\downarrow$) H 426.37 403.58 373.14 ↓53.23 347.21 334.18 303.38 ↓43.83 273.85 257.24 225.28 ↓48.57
M 446.28 436.94 423.75 ↓22.53 385.72 389.24 369.15 ↓16.57 326.92 321.16 302.53 ↓24.39
L 503.38 493.27 473.96 ↓29.42 426.33 419.25 404.28 ↓22.05 372.62 367.26 341.34 ↓31.28
VL 584.55 569.34 532.86 ↓51.69 479.04 463.36 433.48 ↓45.56 396.33 392.33 385.86 ↓10.47
FLORES-200 ($\uparrow$)
(Eng–X)
H 3.14 3.82 5.26 ↑2.12 15.24 16.07 18.47 ↑3.23 23.45 24.33 26.33 ↑2.88
M 2.75 2.94 3.89 ↑1.14 12.83 13.46 15.49 ↑2.66 19.36 20.21 21.62 ↑2.26
L 2.27 2.41 3.14 ↑0.87 8.94 9.28 10.25 ↑1.31 16.61 17.24 18.36 ↑1.75
VL 1.85 2.05 2.35 ↑0.50 6.33 7.25 9.05 ↑2.72 12.51 13.16 14.77 ↑2.26
FLORES-200 ($\uparrow$)
(X–Eng)
H 3.94 3.98 4.26 ↑0.32 16.31 16.92 18.84 ↑2.53 23.86 24.13 26.94 ↑3.08
M 3.52 3.66 3.80 ↑0.28 13.65 14.05 16.27 ↑2.62 20.45 20.36 22.53 ↑2.17
L 3.05 3.12 3.24 ↑0.19 9.47 10.22 11.48 ↑2.01 17.67 17.82 18.93 ↑1.26
VL 2.73 2.83 3.14 ↑0.41 7.28 7.81 9.65 ↑2.37 13.25 13.56 15.88 ↑2.63
Analysis by the Categories of Language Resources. Table 3 presents model performance across languages categorized by resource levels (High, Medium, Low, and Very Low). Across all benchmarks and model sizes, DCAD-2000 consistently outperforms Fineweb-2 and New CC. While gains are modest in high-resource languages, improvements are substantial for low- and very low-resource languages—up to +9.15 accuracy on SIB-200 and -53.23 NLL on Glot500-c—highlighting the effectiveness of our cleaning pipeline in enhancing data quality where it is most needed. The BLEU results on FLORES-200 further validate these trends, with notable improvements in both English-to-X and X-to-English translation tasks. These consistent gains across tasks and languages demonstrate that DCAD-2000 enables more balanced multilingual performance and is well-suited for training inclusive, high-quality language models.
6 Conclusion
In this paper, we introduce DCAD-2000, a large-scale multilingual dataset designed to address the increasing demand for high-quality and diverse training data for multilingual LLMs. Our dataset spans 2,282 languages, providing comprehensive coverage across various geographic regions, scripts (159 scripts), and larger coverage of high/medium resource languages (155 languages). To avoid manually setting thresholds during the data cleaning process, we propose a novel framework that reframes data cleaning as an anomaly detection task. This dynamic approach ensures effective identification and removal of anomalous data from noisy datasets. Empirical experiments demonstrate the effectiveness of our proposed data cleaning framework and the high quality of the DCAD-2000 dataset across multiple multilingual benchmarks.
Appendix A Statistical Analysis of Multilingual Datasets
In this section, we explore the statistical characteristics of the dataset through visual analysis, focusing on the distribution of data across different languages and the variations observed across different shards. We highlight the limitations of existing data cleaning methods that rely on fixed thresholds, particularly in the imbalanced data distribution scenarios. Specifically, when there are substantial discrepancies in word count distributions, these threshold-based cleaning methods are prone to errors, which fail to accurately distinguish between high-quality and low-quality data.
(a) Different Languages
(b) Different Sources
(c) Different Shards
Figure 5: Distribution of average word counts across different languages, sources, and shards in the New CC dataset.
Figure 5(a) illustrates the average word count distribution across different languages in the New CC dataset (CC-MAIN-2024-38). We observe substantial variation in the average word count across languages within the same dataset. For instance, some languages exhibit an average word count as high as 4,000, indicating that their texts are generally longer, while others have an average word count ranging from 50 to 100, suggesting that their texts are typically shorter. This imbalanced distribution complicates the application of traditional fixed-threshold data cleaning methods across all languages. For example, setting a word count threshold of 800 (e.g., the median word count) may be suitable for many languages, but it would still misclassify a significant portion of data as low-quality.
Figure 5(b) illustrates the average word count distribution for Chinese across different data sources (MaLA, Fineweb-2, and New CC). We observe significant variation in the word count distribution for the same language across these sources. For example, the average word count for Chinese in the MaLA corpus is 690, while in New CC (CC-MAIN-2024-33), the average word count increases to 1,975. This discrepancy highlights the inadequacy of a single fixed threshold for data from different sources. Applying a uniform threshold could lead to incorrect cleaning of Chinese text from certain data sources, potentially compromising the representativeness and quality of the data. Consequently, it is essential to adopt flexible cleaning strategies tailored to the characteristics of each data source.
Figure 5(c) illustrates the variation in word count for Chinese across different shards in the Fineweb-2 dataset. We observe imbalanced word count distributions between shards, which further complicates the data cleaning process. For instance, some shards contain texts with word counts concentrated between 700 and 1,000, while others have texts primarily between 1,000 and 1,200. This shard-level variation suggests that fixed-threshold cleaning methods may perform inconsistently across different shards, fails to account for the unique characteristics of the data within each shard. Therefore, in the presence of such imbalanced distributions, it is crucial to implement a more flexible data cleaning approach.
Appendix B DCAD-2000 Grouped by Writting Scripts
As mentioned in Section 4, DCAD-2000 contains a total of 159 writing scripts. To provide a comprehensive overview, we list each of these scripts and their corresponding statistical information in Table 4 and Table 5. By presenting this information, we aim to highlight the broad range of writing systems represented by DCAD and emphasize its potential in various linguistic research and applications.
Appendix C Data Cleaning Statistics
In this section, we provide detailed data cleaning statistics (Table 678 and 9) for high-resource, medium-resource, and low-resource languages. For the data cleaning statistics of very low-resource languages, please refer to the open-source data statistics we released.
Appendix D Experimental Setup
In this section, we provide a comprehensive description of our experimental setup, including dataset preparation, model configurations, continued pretraining procedure, data cleaning and anomaly detection pipeline, evaluation metrics, and implementation details.
D.1 Evaluation Benchmarks
FineTask Benchmark.
FineTask is a multilingual, multi‑task benchmark covering nine typologically diverse languages: Chinese, French, Arabic, Russian, Thai, Hindi, Turkish, Swahili, and Telugu. The benchmark spans a wide range of NLP tasks including reading comprehension, common-sense reasoning, natural language understanding, and text generation. FineTask provides four evaluation metrics: Accuracy, Accuracy normalized over character length, Accuracy normalized over token length, and PMI Accuracy. However, according to statistical data, none of these metrics consistently perform well across all languages. Therefore, we chose to use normalized accuracy (norm accuracy) in our evaluation process.
Multilingual Benchmarks.
To analyze performance across varying resource levels, we further evaluate on three established multilingual corpora:
  • SIB‑200 [24]: A suite of topic classification datasets across 205 languages. We use the raw accuracy on held‑out test sets.
  • Glot500‑c [8]: A curated corpus spanning 500 languages for generation and language modeling. We compute negative log‑likelihood (NLL) on held‑out sentences:
    $\mathrm{NLL}=-\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(w_{t}\mid w_{<t}),$ (9)
    where $T$ is the total token count in the evaluation set.
  • FLORES‑200 [25]: A benchmark for low‑resource machine translation covering 200 languages. We translate from English into each target language (Eng-XX) and translate from other languages into English (X-Eng) and evaluate using SacreBLEU with default settings.
D.2 Pre-training and Evaluation Protocol
We perform continued pretraining on three representative decoder-only large language models (LLMs): LLaMA‑3.2‑1B [47], Qwen‑2.5‑7B [48], and Aya‑expanse‑32B [49]. These models are selected to represent a diverse range of open-source models across different parameter scales, allowing us to investigate the effects of the different dataset, data cleaning pipeline across small, medium, and large model sizes. All models are accessed and managed through the HuggingFace Transformers library.
Given the limitations of computational resources, we refrain from full-parameter finetuning. Instead, we adopt Low-Rank Adaptation (LoRA; 53), a parameter-efficient fine-tuning technique that introduces trainable low-rank matrices into each transformer layer, substantially reducing the number of trainable parameters while maintaining competitive performance.
Our training pipeline closely follows the setup described in the LightEval repository [^16] https://github.com/huggingface/lighteval, a lightweight evaluation and fine-tuning framework developed by HuggingFace. This ensures reproducibility and consistency with widely adopted community practices. Experiments are conducted on NVIDIA A100 GPU with 80GB memory, which provides sufficient memory bandwidth and compute capability to support batch-level parallelism and efficient LoRA-based fine-tuning. All hyperparameters and task-specific configurations are aligned with those used in the Fineweb-2 benchmark, ensuring comparability with previous work and consistent evaluation conditions.
D.3 Statistical Anomaly Detection
We provide a detailed comparison of the anomaly detection algorithms evaluated for data cleaning in DCAD-2000. The methods are selected based on their popularity, conceptual diversity, and availability in scikit-learn [^17] https://scikit-learn.org. All experiments are conducted using the same eight-dimensional feature vectors described in Section 3.
Isolation Forest (ISO_Forest)
[46] is an ensemble-based method that isolates anomalies instead of profiling normal data points. It constructs random binary trees by recursively selecting features and split values, and then uses the path length of each data point across the trees to assess anomaly scores. Shorter paths indicate higher likelihood of being an outlier. ISO_Forest is well-suited to high-dimensional and noisy data and requires minimal hyperparameter tuning. Its main drawback is higher computational cost relative to simpler methods, though it scales well with the number of samples.
One-Class SVM (OC_SVM)
[50] is a kernel-based method that attempts to separate the data from the origin in a transformed feature space. It is sensitive to the choice of kernel function (e.g., RBF, linear) and associated parameters (e.g., gamma, nu). OC_SVM can be effective in capturing complex boundaries, but it often suffers from scalability issues and requires careful parameter tuning, especially in high-dimensional multilingual settings like DCAD-2000.
Local Outlier Factor (LOF)
[51] is a density-based method that identifies anomalies based on local density deviation. It compares the local density of a data point with that of its neighbors. Points that have substantially lower density than their neighbors are considered outliers. The performance of LOF depends heavily on the number of neighbors chosen and tends to degrade in high-dimensional spaces due to the curse of dimensionality. It is also computationally expensive for large datasets.
K-Means
[52] is a clustering algorithm typically used for unsupervised partitioning of data. For anomaly detection, it is repurposed by measuring the distance of points from their assigned cluster centroids—points that are far from any centroid can be considered anomalous. K-Means is computationally efficient and easy to implement but lacks sensitivity to local structures and does not inherently model outliers. Its effectiveness depends on a suitable choice of the number of clusters.
Appendix E Limitations
This work has the following limitations: (i) Although the proportion of high/medium/low resource languages in DCAD-2000 has greatly increased compare to existing multilingual datasets, a significant portion of the languages are still very low resource languages. Future work will explore to collect data for extremely low-resource languages through other modalities (e.g., images) through technologies like OCR. (ii) We evaluate the new data cleaning framework only on four classical anomaly detection algorithms; however, since the framework is algorithm-independent, it should also be effective with other anomaly detection algorithms. (iii) For language identification, we use GlotLID [21], a FastText-based model whose limitations in handling massive multilinguality have been discussed in previous works [54]. However, since the data cleaning pipeline is language-agnostic, other language identification models can also be employed.
Appendix F Ethics Statement
Our dataset integrates existing multilingual datasets, such as MaLA [11] and Fineweb-2 [14], and includes newly extracted data from Common Crawl, providing large-scale and high-quality training corpora to support the training of multilingual large language models (LLMs). Additionally, we propose a novel data cleaning method to filter out potentially toxic documents, reducing potential ethical concerns. However, performing fine-grained analysis on such a vast dataset (46.72TB) remains a significant challenge. To address this, we released the dataset for the community to explore and research extensively. Furthermore, since our dataset is derived from open-source datasets, we adhere to the open-source policies of these datasets to promote future research in multilingual LLMs, while mitigating potential ethical risks. Therefore, we believe our dataset does not pose greater societal risks than existing multilingual datasets.
Table 4: Statisticals grouped by writing scripts (part I). Comparison of language count, document count, token count, disk size, and sources before and after data cleaning in DCAD-2000.
Script #Langs Documents Tokens Disk Size Source
keep remove total keep remove total keep remove total
Latn 1830 5.50B 439.42M 5.93B 4.29T 327.39B 4.61T 21.13TB 4.91TB 26.12TB Fineweb-2, Fineweb, MaLA, New CC
Cyrl 91 1.11B 85.84M 1.19B 1.26T 98.88B 1.36T 9.40TB 2.43TB 11.83TB Fineweb-2, MaLA, New CC
Hani 12 715.15M 71.29M 786.45M 746.48B 73.89B 820.36B 2.90TB 1.60TB 4.50TB Fineweb-2, MaLA, New CC
Jpan 1 491.47M 42.47M 533.93M 278.14B 22.81B 300.95B 2.00TB 504.87GB 2.50TB Fineweb-2, New CC
Arab 60 198.64M 20.36M 219.03M 122.36B 13.12B 135.48B 1.03TB 290.11GB 1.31TB Fineweb-2, MaLA, New CC
Hang 1 79.22M 6.16M 85.38M 59.07B 4.62B 63.69B 336.56GB 66.70GB 403.26GB Fineweb-2, MaLA, New CC
Grek 4 69.14M 5.90M 75.04M 58.45B 5.15B 63.60B 432.64GB 120.10GB 552.76GB Fineweb-2, MaLA, New CC
Deva 48 60.09M 5.79M 65.87M 30.63B 2.56B 33.19B 342.83GB 72.51GB 415.37GB Fineweb-2, MaLA, New CC
Thai 11 55.73M 4.34M 60.06M 46.36B 3.60B 49.96B 526.40GB 110.69GB 637.11GB Fineweb-2, MaLA, New CC
Mlym 6 39.16M 3.89M 43.05M 7.00B 559.61M 7.56B 94.53GB 18.86GB 113.40GB Fineweb-2, MaLA, New CC
Gujr 2 38.91M 4.55M 43.46M 5.07B 461.70M 5.53B 60.22GB 13.54GB 73.76GB Fineweb-2, MaLA, New CC
Knda 2 34.20M 2.70M 36.90M 4.76B 359.45M 5.12B 68.85GB 11.14GB 79.99GB Fineweb-2, MaLA, New CC
Hebr 6 26.99M 1.83M 28.82M 21.15B 1.38B 22.53B 152.34GB 30.80GB 183.18GB Fineweb-2, MaLA, New CC
Taml 2 26.65M 2.92M 29.56M 5.88B 461.60M 6.35B 80.38GB 19.44GB 99.82GB Fineweb-2, MaLA, New CC
Guru 2 24.04M 3.16M 27.21M 2.27B 227.69M 2.50B 26.71GB 8.65GB 35.36GB Fineweb-2, MaLA, New CC
Beng 6 21.91M 1.51M 23.42M 12.67B 875.46M 13.54B 148.42GB 31.39GB 179.83GB Fineweb-2, MaLA, New CC
Geor 3 20.56M 1.36M 21.92M 6.19B 419.61M 6.61B 83.04GB 15.75GB 98.81GB Fineweb-2, MaLA, New CC
Armn 4 17.24M 1.46M 18.70M 4.74B 407.38M 5.15B 42.47GB 11.43GB 53.93GB Fineweb-2, MaLA, New CC
Telu 4 9.93M 821.21K 10.75M 3.91B 295.72M 4.20B 48.22GB 9.65GB 57.87GB Fineweb-2, MaLA, New CC
Sinh 1 9.91M 1.12M 11.03M 2.93B 251.40M 3.18B 32.73GB 7.64GB 40.37GB Fineweb-2, MaLA, New CC
Orya 6 6.57M 616.98K 7.18M 464.57M 37.89M 502.46M 9.79GB 2.20GB 12.01GB Fineweb-2, MaLA
Ethi 13 6.41M 429.99K 6.85M 1.38B 91.75M 1.46B 12.66GB 2.92GB 15.59GB Fineweb-2, MaLA, New CC
Mymr 9 6.04M 479.44K 6.52M 5.30B 406.67M 5.72B 40.57GB 7.83GB 48.39GB Fineweb-2, MaLA, New CC
Kana 1 5.83M 1.11M 6.94M 1.13B 219.26M 1.35B 16.90GB 14.33GB 31.23GB Fineweb-2, New CC
Khmr 7 4.96M 380.38K 5.34M 2.24B 160.29M 2.40B 30.95GB 4.99GB 35.95GB Fineweb-2, MaLA, New CC
Bamu 1 4.71M 1.00M 5.71M 199.46M 42.49M 241.95M 79.67GB 19.47GB 99.14GB Fineweb-2, New CC
Copt 2 4.40M 361.99K 4.76M 219.04M 18.03M 237.09M 8.97GB 864.17MB 9.84GB Fineweb-2, New CC
Tang 1 3.94M 741.81K 4.68M 209.68M 39.47M 249.15M 22.70GB 7.67GB 30.36GB Fineweb-2, New CC
Xsux 1 3.90M 694.59K 4.59M 276.93M 49.35M 326.28M 13.84GB 9.74GB 23.58GB Fineweb-2, New CC
Laoo 5 3.46M 470.52K 3.92M 840.28M 87.36M 927.65M 11.85GB 3.95GB 15.80GB Fineweb-2, MaLA, New CC
Yiii 1 3.39M 417.38K 3.81M 232.88M 28.68M 261.56M 25.82GB 6.24GB 32.05GB Fineweb-2, New CC
Hira 1 2.78M 579.38K 3.36M 361.77M 75.28M 437.05M 4.87GB 4.04GB 8.91GB Fineweb-2, New CC
Thaa 2 2.51M 301.28K 2.82M 425.90M 45.08M 470.98M 4.75GB 1.28GB 6.02GB Fineweb-2, MaLA, New CC
Kits 1 1.86M 315.45K 2.17M 269.54M 45.75M 315.29M 12.47GB 17.12GB 29.58GB Fineweb-2, New CC
Hluw 1 1.71M 374.92K 2.09M 70.77M 15.47M 86.25M 3.19GB 3.45GB 6.64GB Fineweb-2, New CC
Japn 1 1.60M 177.40K 1.78M 148.77M 17.99M 166.76M 6.05GB 2.16GB 8.21GB MaLA
Shrd 1 1.41M 216.59K 1.62M 130.80M 20.13M 150.93M 6.06GB 2.35GB 8.40GB Fineweb-2, New CC
Lina 1 1.37M 271.63K 1.64M 130.39M 25.87M 156.26M 6.97GB 3.85GB 10.82GB Fineweb-2, New CC
Samr 1 1.35M 158.99K 1.51M 64.06M 7.54M 71.59M 4.30GB 1.72GB 6.02GB Fineweb-2, New CC
Cans 12 1.24M 248.84K 1.49M 109.29M 21.66M 130.96M 3.55GB 2.78GB 6.33GB Fineweb-2, MaLA
Syrc 4 1.12M 116.18K 1.23M 44.70M 4.75M 49.44M 20.70GB 4.35GB 25.04GB Fineweb-2, MaLA
Adlm 1 1.12M 194.29K 1.32M 43.63M 7.55M 51.18M 1.10GB 853.95MB 1.95GB Fineweb-2, New CC
Egyp 1 1.12M 190.50K 1.31M 97.41M 16.58M 113.99M 2.54GB 3.52GB 6.05GB Fineweb-2, New CC
Mend 1 1.03M 293.72K 1.32M 16.58M 4.75M 21.33M 893.39MB 2.06GB 2.95GB Fineweb-2, New CC
Linb 1 735.07K 107.67K 842.75K 52.97M 7.76M 60.73M 6.30GB 997.90MB 7.30GB Fineweb-2, New CC
Brai 1 590.10K 125.33K 715.43K 57.85M 12.29M 70.13M 1.94GB 1.30GB 3.24GB Fineweb-2, New CC
Sgnw 1 567.29K 106.45K 673.74K 37.34M 7.01M 44.34M 1.40GB 1.11GB 2.50GB Fineweb-2, New CC
Tibt 4 544.99K 70.33K 615.32K 288.24M 33.57M 321.81M 4.50GB 1.53GB 6.09GB Fineweb-2, MaLA, New CC
Hung 1 520.10K 155.23K 675.33K 42.34M 12.64M 54.98M 1.94GB 2.32GB 4.25GB Fineweb-2, New CC
Mong 3 435.35K 61.47K 496.83K 119.66M 16.95M 136.62M 1.97GB 1.04GB 3.03GB Fineweb-2, MaLA
Bali 1 422.49K 77.08K 499.57K 39.62M 7.23M 46.84M 1.19GB 662.91MB 1.85GB Fineweb-2, New CC
Nshu 1 419.71K 89.40K 509.11K 38.53M 8.21M 46.74M 993.06MB 1.28GB 2.27GB Fineweb-2, New CC
Modi 1 386.82K 67.33K 454.15K 52.58M 9.15M 61.73M 16.45GB 7.42GB 23.87GB Fineweb-2, New CC
Lana 1 377.58K 110.80K 488.38K 47.55M 13.95M 61.50M 688.16MB 2.05GB 2.74GB Fineweb-2, New CC
Saur 1 315.78K 73.82K 389.60K 15.26M 3.57M 18.83M 398.55MB 489.07MB 887.62MB Fineweb-2, New CC
Dupl 1 258.90K 53.06K 311.96K 14.14M 2.90M 17.04M 752.58MB 502.95MB 1.26GB Fineweb-2, New CC
Runr 2 252.18K 39.00K 291.19K 154.68M 23.92M 178.61M 1.25GB 3.28GB 4.52GB Fineweb-2, MaLA
Vaii 1 243.47K 93.27K 336.73K 71.28M 27.31M 98.59M 513.30MB 1.88GB 2.39GB Fineweb-2, New CC
Glag 1 237.68K 72.07K 309.75K 20.38M 6.18M 26.56M 476.61MB 951.96MB 1.43GB Fineweb-2, New CC
Dsrt 1 198.00K 37.90K 235.90K 4.47M 855.49K 5.32M 248.83MB 562.92MB 811.75MB Fineweb-2, New CC
Mroo 1 186.14K 22.85K 208.99K 6.42M 788.69K 7.21M 2.43GB 335.38MB 2.77GB Fineweb-2, New CC
Bopo 1 181.71K 24.45K 206.16K 30.63M 4.12M 34.75M 3.45GB 890.68MB 4.35GB Fineweb-2, New CC
Mtei 2 175.69K 20.34K 196.03K 49.11M 5.76M 54.87M 805.36MB 574.03MB 1.38GB Fineweb-2, MaLA
Khar 1 153.37K 40.04K 193.41K 6.75M 1.76M 8.52M 250.30MB 182.38MB 432.67MB Fineweb-2, New CC
Brah 1 138.03K 22.72K 160.75K 7.85M 1.29M 9.15M 273.71MB 243.75MB 517.47MB Fineweb-2, New CC
Bhks 1 131.90K 27.03K 158.93K 3.93M 805.58K 4.74M 190.96MB 154.63MB 345.59MB Fineweb-2, New CC
Hmnp 1 118.87K 12.33K 131.20K 6.83M 708.37K 7.54M 436.28MB 151.81MB 588.09MB Fineweb-2, New CC
Phag 1 107.75K 17.58K 125.34K 3.41M 556.36K 3.97M 141.68MB 93.31MB 234.99MB Fineweb-2, New CC
Merc 1 107.52K 38.04K 145.56K 7.61M 2.69M 10.30M 215.43MB 472.23MB 687.66MB Fineweb-2, New CC
Kali 2 105.87K 24.33K 130.20K 1.39M 319.46K 1.71M 105.24MB 91.45MB 196.70MB Fineweb-2, New CC
Plrd 1 104.31K 21.07K 125.38K 5.47M 1.10M 6.57M 214.53MB 225.25MB 439.77MB Fineweb-2, New CC
Lisu 2 101.48K 20.06K 121.53K 24.00M 4.74M 28.74M 204.24MB 527.21MB 731.45MB Fineweb-2, New CC
Hmng 1 101.02K 23.34K 124.36K 5.37M 1.24M 6.61M 153.20MB 196.99MB 350.19MB Fineweb-2, New CC
Nkoo 2 98.77K 25.89K 124.65K 4.91M 1.07M 5.98M 2.13GB 233.87MB 2.36GB Fineweb-2, MaLA
Gran 1 97.96K 21.57K 119.53K 3.57M 785.93K 4.36M 135.27MB 243.90MB 379.18MB Fineweb-2, New CC
Gonm 1 94.82K 16.28K 111.10K 2.83M 486.36K 3.32M 106.89MB 142.16MB 249.05MB Fineweb-2, New CC
Cher 2 94.19K 25.99K 120.19K 9.12M 2.45M 11.57M 245.29MB 689.18MB 934.47MB Fineweb-2, MaLA
Tnsa 1 89.55K 17.93K 107.48K 3.28M 656.33K 3.93M 98.49MB 204.04MB 302.53MB Fineweb-2, New CC
Table 5: Statisticals grouped by writing scripts (part II). Comparison of language count, document count, token count, disk size, and sources before and after data cleaning in DCAD-2000.
Script #Langs Documents Tokens Disk Size Source
keep remove total keep remove total keep remove total
Cprt 1 88.19K 14.11K 102.30K 7.87M 1.26M 9.13M 142.36MB 85.91MB 228.27MB Fineweb-2, New CC
Cari 1 77.73K 18.09K 95.82K 1.73M 401.78K 2.13M 89.37MB 76.01MB 165.38MB Fineweb-2, New CC
Diak 1 68.42K 22.40K 90.82K 2.87M 938.52K 3.81M 58.40MB 94.36MB 152.76MB Fineweb-2, New CC
Marc 1 67.80K 11.89K 79.69K 2.34M 410.50K 2.75M 66.51MB 95.34MB 161.85MB Fineweb-2, New CC
Mani 1 65.94K 9.56K 75.50K 6.27M 908.84K 7.17M 128.39MB 140.35MB 268.75MB Fineweb-2, New CC
Talu 2 65.77K 11.95K 77.72K 1.27M 231.55K 1.50M 78.51MB 62.21MB 140.72MB Fineweb-2, MaLA
Vith 1 65.14K 12.13K 77.28K 2.49M 464.49K 2.96M 124.41MB 95.26MB 219.66MB Fineweb-2, New CC
Nagm 1 63.57K 11.94K 75.51K 1.03M 193.45K 1.22M 58.20MB 73.87MB 132.08MB Fineweb-2, New CC
Ahom 1 60.21K 9.69K 69.90K 2.34M 376.34K 2.72M 127.53MB 70.68MB 198.21MB Fineweb-2, New CC
Java 1 58.52K 13.32K 71.84K 2.18M 496.30K 2.68M 66.55MB 116.13MB 182.68MB Fineweb-2, New CC
Palm 1 48.99K 5.32K 54.32K 424.13K 46.09K 470.22K 39.41MB 43.82MB 83.23MB Fineweb-2, New CC
Wara 1 46.80K 9.12K 55.92K 1.47M 286.76K 1.76M 58.48MB 52.76MB 111.24MB Fineweb-2, New CC
Olck 2 45.80K 4.06K 49.86K 6.69M 492.54K 7.19M 86.16MB 38.55MB 124.71MB Fineweb-2, MaLA
Khoj 1 39.85K 5.23K 45.09K 892.46K 117.20K 1.01M 43.07MB 40.20MB 83.27MB Fineweb-2, New CC
Rohg 1 35.21K 5.32K 40.53K 534.34K 80.72K 615.05K 36.76MB 41.06MB 77.82MB Fineweb-2, New CC
Sidd 1 34.75K 8.41K 43.16K 3.03M 732.80K 3.76M 46.06MB 93.44MB 139.51MB Fineweb-2, New CC
Yezi 1 33.92K 3.35K 37.27K 96.61K 9.53K 106.13K 29.36MB 14.31MB 43.67MB Fineweb-2, New CC
Ougr 1 32.34K 6.13K 38.47K 442.16K 83.82K 525.98K 31.03MB 37.95MB 68.98MB Fineweb-2, New CC
Avst 1 32.16K 6.62K 38.78K 1.75M 360.09K 2.11M 51.64MB 53.81MB 105.46MB Fineweb-2, New CC
Ital 1 32.06K 5.06K 37.12K 519.27K 81.93K 601.19K 34.30MB 29.24MB 63.53MB Fineweb-2, New CC
Wcho 1 31.94K 6.51K 38.45K 1.48M 301.04K 1.78M 58.25MB 74.54MB 132.79MB Fineweb-2, New CC
Kthi 1 31.07K 5.44K 36.51K 763.52K 133.75K 897.27K 30.79MB 35.73MB 66.52MB Fineweb-2, New CC
Tavt 1 30.95K 3.63K 34.57K 670.82K 78.65K 749.47K 29.30MB 14.97MB 44.26MB Fineweb-2, New CC
Takr 1 30.70K 5.29K 35.99K 1.73M 298.02K 2.03M 30.89MB 45.59MB 76.48MB Fineweb-2, New CC
Tfng 4 29.84K 3.34K 33.18K 1.42M 148.55K 1.57M 35.12MB 24.87MB 59.99MB Fineweb-2, New CC
Tale 1 26.17K 2.80K 28.98K 220.84K 23.64K 244.48K 23.80MB 16.84MB 40.64MB Fineweb-2, New CC
Elba 1 24.86K 4.61K 29.48K 394.51K 73.22K 467.73K 24.19MB 19.19MB 43.38MB Fineweb-2, New CC
Zanb 1 24.46K 4.76K 29.21K 327.39K 63.68K 391.07K 26.07MB 40.03MB 66.10MB Fineweb-2, New CC
Sogo 1 22.29K 3.88K 26.16K 146.13K 25.41K 171.54K 17.82MB 20.07MB 37.89MB Fineweb-2, New CC
Soyo 1 22.21K 4.91K 27.12K 598.89K 132.47K 731.36K 25.04MB 36.77MB 61.81MB Fineweb-2, New CC
Dogr 1 21.29K 3.82K 25.11K 1.28M 229.94K 1.51M 29.94MB 23.89MB 53.84MB Fineweb-2, New CC
Kawi 1 20.28K 4.10K 24.38K 396.57K 80.26K 476.83K 20.90MB 24.30MB 45.20MB Fineweb-2, New CC
Phli 1 19.16K 2.88K 22.04K 41.16K 6.19K 47.35K 17.52MB 7.60MB 25.13MB Fineweb-2, New CC
Cham 1 17.92K 3.60K 21.52K 762.24K 153.32K 915.57K 21.12MB 39.91MB 61.03MB Fineweb-2, New CC
Nbat 1 17.61K 3.19K 20.80K 280.13K 50.76K 330.89K 18.90MB 15.97MB 34.87MB Fineweb-2, New CC
Nand 1 17.39K 3.36K 20.75K 307.12K 59.32K 366.44K 17.76MB 19.20MB 36.96MB Fineweb-2, New CC
Osma 1 16.98K 2.59K 19.57K 495.54K 75.61K 571.15K 19.16MB 15.11MB 34.27MB Fineweb-2, New CC
Sind 1 14.81K 4.24K 19.05K 315.61K 90.31K 405.93K 21.16MB 18.70MB 39.86MB Fineweb-2, New CC
Sogd 1 14.52K 2.73K 17.24K 307.50K 57.79K 365.30K 14.67MB 9.73MB 24.40MB Fineweb-2, New CC
Pauc 1 13.23K 4.28K 17.50K 1.88M 609.43K 2.49M 13.65MB 33.03MB 46.67MB Fineweb-2, New CC
Sylo 1 12.42K 2.88K 15.29K 922.71K 213.86K 1.14M 22.76MB 22.23MB 44.99MB Fineweb-2, New CC
Goth 2 11.84K 1.24K 13.08K 191.30K 19.67K 210.97K 11.59MB 3.62MB 15.22MB Fineweb-2, MaLA
Rjng 1 10.30K 2.36K 12.65K 595.51K 136.27K 731.78K 9.43MB 15.02MB 24.45MB Fineweb-2, New CC
Chrs 1 10.24K 1.26K 11.50K 45.98K 5.66K 51.64K 8.22MB 5.45MB 13.67MB Fineweb-2, New CC
Phlp 1 9.08K 2.03K 11.11K 31.62K 7.06K 38.69K 8.35MB 5.61MB 13.96MB Fineweb-2, New CC
Mand 1 8.73K 1.49K 10.21K 82.87K 14.11K 96.98K 9.07MB 5.24MB 14.31MB Fineweb-2, New CC
Tglg 1 8.58K 1.88K 10.46K 638.75K 140.15K 778.89K 11.22MB 10.89MB 22.11MB Fineweb-2, New CC
Shaw 1 8.41K 1.28K 9.69K 915.43K 139.72K 1.06M 13.65MB 12.62MB 26.27MB Fineweb-2, New CC
Hatr 1 7.44K 1.63K 9.07K 371.48K 81.61K 453.09K 10.15MB 13.53MB 23.68MB Fineweb-2, New CC
Bugi 2 7.03K 1.33K 8.36K 95.81K 18.11K 113.91K 6.90MB 6.18MB 13.09MB Fineweb-2, MaLA
Tagb 1 6.58K 1.14K 7.72K 30.92K 5.37K 36.30K 5.84MB 2.33MB 8.17MB Fineweb-2, New CC
Prti 1 6.05K 1.09K 7.15K 225.93K 40.79K 266.72K 7.31MB 4.57MB 11.89MB Fineweb-2, New CC
Narb 1 5.22K 835 6.06K 56.09K 8.97K 65.06K 6.01MB 7.12MB 13.13MB Fineweb-2, New CC
Sarb 1 4.99K 874 5.86K 170.46K 29.86K 200.31K 6.93MB 15.95MB 22.87MB Fineweb-2, New CC
Ugar 1 4.85K 653 5.50K 133.05K 17.92K 150.97K 4.03MB 2.47MB 6.50MB Fineweb-2, New CC
Lydi 1 4.59K 1.03K 5.62K 28.08M 6.29M 34.37M 77.22MB 70.99MB 148.21MB Fineweb-2, New CC
Buhd 1 3.16K 448 3.61K 7.77K 1.10K 8.87K 2.73MB 623.88KB 3.35MB Fineweb-2, New CC
Perm 1 2.87K 630 3.50K 19.17K 4.20K 23.37K 2.58MB 1.36MB 3.94MB Fineweb-2, New CC
Elym 1 1.66K 496 2.16K 61.25K 18.28K 79.53K 1.88MB 7.52MB 9.40MB Fineweb-2, New CC
Limb 1 59 15 74 32.32K 8.22K 40.53K 754.75KB 229.80KB 984.54KB Fineweb-2, New CC
Table 6: Data Cleaning Statistics (part I): Comparison of document count, token count, disk size, and sources before and after data cleaning in DCAD-2000.
Lang Code Documents Tokens Disk Size Source
keep remove total keep remove total keep remove total
eng_Latn 1.31B 101.08M 1.41B 1.21T 93.23B 1.30T 5.66TB 1.49TB 7.15TB Fineweb, MaLA, New CC
rus_Cyrl 858.53M 67.21M 925.74M 1.14T 90.18B 1.23T 8.40TB 2.22TB 10.62TB Fineweb-2, MaLA, New CC
cmn_Hani 713.97M 71.19M 785.16M 745.88B 73.84B 819.71B 2.90TB 1.60TB 4.50TB Fineweb-2, New CC
deu_Latn 668.62M 53.65M 722.27M 632.32B 51.11B 683.44B 2.85TB 664.79GB 3.52TB Fineweb-2, MaLA, New CC
spa_Latn 604.45M 43.33M 647.79M 483.75B 34.79B 518.54B 2.54TB 498.55GB 3.03TB Fineweb-2, MaLA, New CC
fra_Latn 513.53M 40.32M 553.85M 430.86B 33.77B 464.64B 2.15TB 491.23GB 2.64TB Fineweb-2, MaLA, New CC
jpn_Jpan 491.47M 42.47M 533.93M 278.14B 22.81B 300.95B 2.00TB 504.87GB 2.50TB Fineweb-2, New CC
ita_Latn 311.42M 25.50M 336.93M 250.12B 20.69B 270.81B 1.29TB 292.75GB 1.59TB Fineweb-2, MaLA, New CC
por_Latn 271.48M 18.67M 290.15M 204.64B 14.12B 218.77B 1.07TB 225.83GB 1.30TB Fineweb-2, MaLA, New CC
pol_Latn 223.21M 15.38M 238.59M 180.34B 12.59B 192.93B 910.55GB 184.59GB 1.10TB Fineweb-2, MaLA, New CC
nld_Latn 219.16M 14.03M 233.19M 146.16B 9.38B 155.54B 739.62GB 159.01GB 898.63GB Fineweb-2, MaLA, New CC
ind_Latn 156.92M 16.21M 173.12M 60.97B 5.15B 66.11B 406.86GB 64.84GB 471.70GB Fineweb-2, MaLA
tur_Latn 143.31M 9.98M 153.30M 118.40B 8.21B 126.61B 618.87GB 145.39GB 764.26GB Fineweb-2, MaLA, New CC
vie_Latn 87.77M 6.19M 93.96M 110.11B 7.71B 117.82B 570.86GB 116.19GB 687.05GB Fineweb-2, MaLA, New CC
fas_Arab 82.80M 9.49M 92.29M 67.58B 7.91B 75.49B 521.39GB 121.46GB 642.85GB Fineweb-2, MaLA, New CC
kor_Hang 79.22M 6.16M 85.38M 59.07B 4.62B 63.69B 336.56GB 66.70GB 403.26GB Fineweb-2, MaLA, New CC
swe_Latn 77.32M 5.08M 82.40M 59.21B 3.92B 63.13B 269.37GB 73.25GB 342.62GB Fineweb-2, MaLA, New CC
hun_Latn 70.79M 5.18M 75.97M 65.62B 4.86B 70.48B 319.58GB 87.97GB 407.55GB Fineweb-2, MaLA, New CC
ukr_Cyrl 67.87M 4.31M 72.18M 53.79B 3.41B 57.20B 428.74GB 82.51GB 511.25GB Fineweb-2, MaLA, New CC
ell_Grek 67.48M 5.67M 73.15M 57.63B 5.03B 62.66B 425.03GB 112.67GB 537.71GB Fineweb-2, MaLA, New CC
tha_Thai 55.47M 4.29M 59.76M 46.31B 3.59B 49.90B 525.54GB 110.37GB 635.91GB Fineweb-2, MaLA, New CC
arb_Arab 53.70M 4.06M 57.76M 25.21B 1.92B 27.13B 278.77GB 72.25GB 351.02GB Fineweb-2, MaLA
aze_Latn 51.38M 6.44M 57.82M 3.30B 392.00M 3.70B 41.90GB 10.70GB 52.60GB MaLA
slv_Latn 50.41M 4.05M 54.46M 11.66B 836.48M 12.50B 69.22GB 12.64GB 81.87GB Fineweb-2, MaLA, New CC
cat_Latn 48.83M 3.78M 52.61M 16.49B 1.13B 17.62B 96.97GB 14.24GB 111.21GB Fineweb-2, MaLA, New CC
fin_Latn 47.80M 4.09M 51.89M 43.43B 3.75B 47.19B 202.14GB 57.62GB 259.76GB Fineweb-2, MaLA, New CC
ces_Latn 47.54M 3.21M 50.74M 42.20B 2.84B 45.04B 195.62GB 48.74GB 244.36GB MaLA, New CC
hbs_Latn 42.98M 8.05M 51.04M 1.53B 287.34M 1.82B 22.41GB 6.41GB 28.82GB MaLA
fil_Latn 40.15M 6.32M 46.47M 3.47B 477.70M 3.94B 31.22GB 9.20GB 40.42GB Fineweb-2, MaLA
mal_Mlym 39.10M 3.88M 42.98M 7.00B 558.83M 7.56B 94.47GB 18.30GB 112.78GB Fineweb-2, MaLA, New CC
nob_Latn 38.88M 4.33M 43.21M 24.13B 2.81B 26.94B 139.85GB 66.29GB 206.15GB Fineweb-2, MaLA
guj_Gujr 38.82M 4.54M 43.36M 5.07B 461.54M 5.53B 60.08GB 13.49GB 73.57GB Fineweb-2, MaLA, New CC
bul_Cyrl 37.11M 2.56M 39.67M 32.29B 2.23B 34.51B 245.84GB 55.86GB 301.69GB Fineweb-2, MaLA, New CC
kan_Knda 34.20M 2.70M 36.90M 4.76B 359.21M 5.12B 68.82GB 11.13GB 79.95GB Fineweb-2, MaLA, New CC
hin_Deva 29.15M 2.47M 31.62M 22.08B 1.81B 23.89B 219.46GB 46.45GB 265.91GB Fineweb-2, MaLA, New CC
tam_Taml 26.55M 2.90M 29.45M 5.88B 460.75M 6.34B 80.26GB 19.29GB 99.55GB Fineweb-2, MaLA, New CC
kaz_Cyrl 25.78M 1.67M 27.45M 6.37B 432.67M 6.80B 64.36GB 12.99GB 77.35GB Fineweb-2, MaLA, New CC
heb_Hebr 25.24M 1.61M 26.85M 20.74B 1.33B 22.07B 147.85GB 28.75GB 176.60GB Fineweb-2, MaLA, New CC
ara_Arab 25.14M 3.24M 28.39M 17.21B 2.23B 19.44B 152.73GB 71.93GB 224.66GB MaLA, New CC
srp_Cyrl 25.13M 1.75M 26.88M 6.91B 496.07M 7.41B 60.34GB 8.50GB 68.84GB Fineweb-2, MaLA, New CC
est_Latn 24.18M 2.86M 27.04M 2.89B 294.20M 3.18B 26.17GB 8.91GB 35.08GB MaLA, New CC
sqi_Latn 24.16M 3.25M 27.41M 2.38B 237.81M 2.61B 21.08GB 5.03GB 26.11GB MaLA, New CC
isl_Latn 24.06M 2.23M 26.29M 6.32B 561.74M 6.89B 34.88GB 9.09GB 43.97GB Fineweb-2, MaLA, New CC
pan_Guru 24.02M 3.16M 27.19M 2.27B 227.60M 2.50B 26.69GB 8.59GB 35.28GB MaLA, New CC
mlt_Latn 23.37M 2.08M 25.45M 3.24B 322.80M 3.56B 16.40GB 4.96GB 21.36GB Fineweb-2, MaLA, New CC
mkd_Cyrl 22.61M 1.89M 24.50M 5.29B 396.98M 5.68B 51.37GB 7.08GB 58.45GB Fineweb-2, MaLA, New CC
bos_Latn 21.62M 1.71M 23.33M 11.01B 831.59M 11.84B 59.71GB 10.67GB 70.38GB Fineweb-2, MaLA, New CC
kat_Geor 20.27M 1.30M 21.57M 6.16B 413.36M 6.57B 82.54GB 15.10GB 97.65GB Fineweb-2, MaLA, New CC
lit_Latn 20.09M 1.51M 21.60M 17.47B 1.33B 18.80B 91.29GB 18.30GB 109.59GB Fineweb-2, MaLA, New CC
ben_Beng 19.90M 1.37M 21.28M 12.26B 848.75M 13.11B 143.64GB 30.36GB 174.00GB Fineweb-2, MaLA, New CC
hrv_Latn 19.83M 1.54M 21.37M 15.02B 1.19B 16.21B 76.53GB 16.65GB 93.18GB Fineweb-2, MaLA, New CC
glg_Latn 19.31M 1.58M 20.89M 4.45B 372.72M 4.83B 28.40GB 4.50GB 32.90GB Fineweb-2, MaLA, New CC
ron_Latn 18.28M 1.42M 19.69M 23.42B 1.81B 25.23B 110.94GB 20.14GB 131.08GB MaLA, New CC
ceb_Latn 18.14M 1.82M 19.97M 1.91B 184.52M 2.09B 14.11GB 2.06GB 16.18GB Fineweb-2, MaLA, New CC
hye_Armn 16.93M 1.40M 18.33M 4.65B 392.68M 5.04B 41.29GB 10.76GB 52.05GB Fineweb-2, MaLA, New CC
msa_Latn 16.90M 1.51M 18.40M 12.27B 1.05B 13.32B 67.19GB 34.22GB 101.42GB MaLA, New CC
tgk_Cyrl 16.60M 1.04M 17.64M 3.46B 241.47M 3.70B 29.00GB 5.01GB 34.01GB Fineweb-2, MaLA, New CC
mar_Deva 15.37M 1.35M 16.72M 4.05B 287.28M 4.34B 52.49GB 7.16GB 59.65GB Fineweb-2, MaLA, New CC
bel_Cyrl 15.22M 1.06M 16.29M 5.30B 353.85M 5.65B 45.23GB 6.76GB 51.99GB Fineweb-2, MaLA, New CC
nep_Deva 13.18M 1.74M 14.91M 3.40B 354.95M 3.75B 57.72GB 14.16GB 71.88GB MaLA, New CC
urd_Arab 12.92M 1.28M 14.20M 5.63B 463.49M 6.09B 43.36GB 8.33GB 51.69GB Fineweb-2, MaLA, New CC
slk_Latn 12.79M 850.42K 13.64M 10.71B 712.57M 11.43B 53.49GB 10.01GB 63.50GB MaLA, New CC
mon_Cyrl 11.46M 1.37M 12.83M 2.05B 225.17M 2.27B 25.55GB 7.89GB 33.44GB MaLA, New CC
dan_Latn 11.33M 645.36K 11.98M 8.91B 506.75M 9.42B 42.48GB 9.31GB 51.78GB MaLA, New CC
eus_Latn 10.88M 720.92K 11.60M 2.86B 180.73M 3.04B 18.54GB 2.98GB 21.52GB Fineweb-2, MaLA, New CC
azj_Latn 10.37M 764.57K 11.14M 6.02B 427.97M 6.45B 54.46GB 9.98GB 64.44GB Fineweb-2, MaLA, New CC
swa_Latn 10.32M 1.78M 12.10M 968.63M 131.70M 1.10B 8.88GB 2.59GB 11.47GB MaLA, New CC
als_Latn 9.94M 695.21K 10.64M 7.84B 540.49M 8.38B 22.16GB 3.80GB 25.97GB Fineweb-2, MaLA
sin_Sinh 9.91M 1.12M 11.03M 2.93B 251.40M 3.18B 32.73GB 7.64GB 40.37GB Fineweb-2, MaLA, New CC
lat_Latn 9.86M 968.13K 10.83M 1.67B 209.54M 1.88B 8.93GB 3.35GB 12.27GB Fineweb-2, MaLA, New CC
tel_Telu 9.81M 790.37K 10.60M 3.90B 293.32M 4.19B 47.82GB 9.23GB 57.05GB Fineweb-2, MaLA, New CC
afr_Latn 9.38M 858.54K 10.24M 3.02B 252.81M 3.27B 16.05GB 3.08GB 19.13GB Fineweb-2, MaLA, New CC
Table 7: Data Cleaning Statistics (part II): Comparison of document count, token count, disk size, and sources before and after data cleaning in DCAD-2000.
Lang Code Documents Tokens Disk Size Source
keep remove total keep remove total keep remove total
ekk_Latn 9.24M 772.47K 10.01M 4.79B 401.83M 5.19B 38.34GB 11.83GB 50.16GB Fineweb-2, MaLA
zsm_Latn 8.67M 795.54K 9.47M 4.22B 365.48M 4.59B 31.54GB 8.93GB 40.48GB Fineweb-2, MaLA
ltz_Latn 8.59M 1.21M 9.79M 1.18B 146.26M 1.33B 6.77GB 1.91GB 8.68GB Fineweb-2, MaLA, New CC
som_Latn 7.47M 716.70K 8.19M 2.20B 193.46M 2.40B 10.27GB 3.34GB 13.61GB Fineweb-2, MaLA, New CC
kir_Cyrl 6.47M 468.94K 6.94M 2.31B 183.29M 2.49B 21.00GB 3.63GB 24.63GB Fineweb-2, MaLA, New CC
cym_Latn 6.47M 515.43K 6.99M 2.01B 141.85M 2.15B 10.29GB 1.99GB 12.28GB Fineweb-2, MaLA, New CC
nor_Latn 6.13M 733.57K 6.87M 1.27B 150.12M 1.42B 8.91GB 2.74GB 11.65GB MaLA, New CC
uzb_Latn 6.07M 715.37K 6.78M 929.54M 98.71M 1.03B 8.76GB 2.73GB 11.49GB MaLA, New CC
und_Kana 5.83M 1.11M 6.94M 1.13B 219.26M 1.35B 16.90GB 14.33GB 31.23GB Fineweb-2, New CC
mya_Mymr 5.80M 449.02K 6.25M 5.28B 404.36M 5.69B 40.05GB 7.53GB 47.57GB Fineweb-2, MaLA, New CC
epo_Latn 5.77M 456.78K 6.23M 2.38B 177.31M 2.56B 12.03GB 2.25GB 14.27GB Fineweb-2, MaLA, New CC
ary_Arab 5.67M 465.36K 6.14M 1.38B 114.17M 1.50B 18.12GB 4.32GB 22.44GB Fineweb-2, MaLA
lvs_Latn 5.51M 382.81K 5.89M 2.74B 185.99M 2.92B 21.58GB 6.85GB 28.43GB Fineweb-2, MaLA
hau_Latn 5.48M 662.28K 6.15M 438.94M 49.38M 488.32M 3.22GB 1.09GB 4.32GB MaLA
gle_Latn 5.47M 428.92K 5.90M 1.65B 134.54M 1.78B 9.41GB 1.55GB 10.96GB Fineweb-2, MaLA, New CC
nno_Latn 5.19M 553.48K 5.75M 1.35B 124.05M 1.48B 7.48GB 1.85GB 9.33GB Fineweb-2, MaLA, New CC
ory_Orya 5.13M 444.55K 5.57M 325.74M 23.33M 349.07M 7.34GB 1.07GB 8.41GB Fineweb-2, MaLA
amh_Ethi 4.86M 302.32K 5.17M 1.21B 77.95M 1.28B 10.27GB 1.56GB 11.83GB Fineweb-2, MaLA, New CC
khm_Khmr 4.74M 344.10K 5.09M 2.23B 158.45M 2.39B 30.49GB 4.58GB 35.08GB Fineweb-2, MaLA, New CC
tat_Cyrl 4.72M 390.38K 5.11M 1.29B 103.35M 1.39B 11.66GB 2.16GB 13.82GB Fineweb-2, MaLA, New CC
und_Bamu 4.71M 1.00M 5.71M 199.46M 42.49M 241.95M 79.67GB 19.47GB 99.14GB Fineweb-2, New CC
und_Copt 4.40M 361.86K 4.76M 218.11M 17.95M 236.07M 8.96GB 860.56MB 9.82GB Fineweb-2, New CC
arz_Arab 4.19M 347.36K 4.54M 794.23M 62.86M 857.09M 6.87GB 1.16GB 8.03GB Fineweb-2, MaLA, New CC
und_Tang 3.94M 741.81K 4.68M 209.68M 39.47M 249.15M 22.70GB 7.67GB 30.36GB Fineweb-2, New CC
und_Xsux 3.90M 694.59K 4.59M 276.93M 49.35M 326.28M 13.84GB 9.74GB 23.58GB Fineweb-2, New CC
lav_Latn 3.76M 347.11K 4.11M 2.12B 196.45M 2.31B 13.96GB 7.36GB 21.32GB MaLA, New CC
pus_Arab 3.71M 493.24K 4.21M 905.77M 106.28M 1.01B 7.66GB 2.36GB 10.02GB MaLA, New CC
hbs_Cyrl 3.47M 463.55K 3.93M 131.15M 17.53M 148.69M 2.47GB 544.73MB 3.02GB MaLA, New CC
war_Latn 3.43M 283.72K 3.71M 137.36M 11.19M 148.55M 1.84GB 161.55MB 2.00GB Fineweb-2, MaLA, New CC
und_Yiii 3.39M 417.38K 3.81M 232.88M 28.68M 261.56M 25.82GB 6.24GB 32.05GB Fineweb-2, New CC
multi_Latn 3.11M 394.01K 3.50M 2.39B 303.45M 2.70B 18.42GB 7.60GB 26.02GB New CC
mlg_Latn 2.85M 437.74K 3.29M 288.34M 41.29M 329.63M 2.74GB 765.89MB 3.51GB MaLA, New CC
und_Hira 2.78M 579.38K 3.36M 361.77M 75.28M 437.05M 4.87GB 4.04GB 8.91GB Fineweb-2, New CC
uzn_Cyrl 2.61M 304.12K 2.91M 396.89M 30.84M 427.73M 6.39GB 1.47GB 7.86GB Fineweb-2, MaLA
hat_Latn 2.58M 226.91K 2.81M 464.18M 41.25M 505.43M 2.60GB 548.40MB 3.15GB Fineweb-2, MaLA, New CC
zul_Latn 2.47M 294.21K 2.76M 333.05M 38.27M 371.33M 2.15GB 642.83MB 2.79GB Fineweb-2, MaLA
kur_Latn 2.41M 327.93K 2.74M 482.02M 51.67M 533.69M 3.40GB 1.04GB 4.44GB MaLA
div_Thaa 2.25M 263.72K 2.52M 418.22M 43.98M 462.20M 4.37GB 1.02GB 5.38GB Fineweb-2, MaLA, New CC
tgl_Latn 2.24M 345.69K 2.59M 369.19M 35.56M 404.74M 2.75GB 669.71MB 3.42GB MaLA, New CC
uzb_Cyrl 2.22M 314.25K 2.54M 194.02M 27.60M 221.61M 2.96GB 1.14GB 4.10GB MaLA
fry_Latn 2.14M 232.49K 2.38M 605.32M 65.90M 671.22M 3.10GB 914.11MB 4.01GB Fineweb-2, MaLA, New CC
sna_Latn 2.14M 181.61K 2.32M 295.33M 24.54M 319.87M 1.84GB 428.76MB 2.27GB Fineweb-2, MaLA
fao_Latn 2.09M 163.66K 2.26M 199.43M 14.19M 213.61M 1.69GB 392.84MB 2.08GB Fineweb-2, MaLA
und_Laoo 2.06M 364.70K 2.42M 212.14M 37.64M 249.78M 4.20GB 2.59GB 6.79GB Fineweb-2, New CC
sun_Latn 1.99M 193.82K 2.19M 275.24M 25.28M 300.53M 1.71GB 543.58MB 2.25GB Fineweb-2, MaLA, New CC
snd_Arab 1.91M 154.84K 2.06M 1.12B 105.00M 1.22B 5.27GB 1.88GB 7.15GB Fineweb-2, MaLA, New CC
und_Cyrl 1.86M 427.20K 2.29M 1.32B 302.88M 1.62B 5.09GB 18.81GB 23.90GB Fineweb-2, New CC
und_Kits 1.86M 315.45K 2.17M 269.54M 45.75M 315.29M 12.47GB 17.12GB 29.58GB Fineweb-2, New CC
bak_Cyrl 1.85M 132.43K 1.99M 401.91M 27.62M 429.53M 3.87GB 733.50MB 4.60GB Fineweb-2, MaLA, New CC
asm_Beng 1.82M 115.52K 1.93M 380.78M 23.67M 404.45M 4.50GB 907.15MB 5.40GB Fineweb-2, MaLA, New CC
cos_Latn 1.79M 274.66K 2.06M 228.06M 35.24M 263.31M 1.10GB 580.00MB 1.68GB MaLA
ckb_Arab 1.78M 177.88K 1.96M 841.60M 76.59M 918.19M 6.48GB 1.52GB 8.00GB Fineweb-2, MaLA, New CC
und_Hluw 1.71M 374.92K 2.09M 70.77M 15.47M 86.25M 3.19GB 3.45GB 6.64GB Fineweb-2, New CC
ast_Latn 1.63M 144.18K 1.77M 213.12M 19.08M 232.20M 1.39GB 385.45MB 1.78GB Fineweb-2, MaLA, New CC
jpn_Japn 1.60M 177.40K 1.78M 148.77M 17.99M 166.76M 6.05GB 2.16GB 8.21GB MaLA
ibo_Latn 1.59M 117.64K 1.71M 233.50M 16.65M 250.14M 1.45GB 446.07MB 1.89GB Fineweb-2, MaLA
und_Grek 1.57M 224.66K 1.79M 755.84M 108.19M 864.02M 6.94GB 7.17GB 14.12GB Fineweb-2, New CC
mri_Latn 1.53M 133.72K 1.67M 354.50M 28.72M 383.22M 1.71GB 472.53MB 2.18GB Fineweb-2, MaLA
ars_Arab 1.53M 108.78K 1.64M 461.05M 32.76M 493.81M 4.88GB 1.85GB 6.73GB Fineweb-2, New CC
anp_Deva 1.44M 140.26K 1.58M 805.49M 78.54M 884.04M 10.69GB 2.12GB 12.81GB Fineweb-2, MaLA
khk_Cyrl 1.44M 128.14K 1.57M 615.04M 54.80M 669.84M 8.17GB 1.83GB 10.00GB Fineweb-2, New CC
und_Shrd 1.41M 216.59K 1.62M 130.80M 20.13M 150.93M 6.06GB 2.35GB 8.40GB Fineweb-2, New CC
lao_Laoo 1.40M 105.80K 1.50M 628.08M 49.71M 677.79M 7.65GB 1.36GB 9.01GB Fineweb-2, MaLA, New CC
und_Lina 1.37M 271.63K 1.64M 130.39M 25.87M 156.26M 6.97GB 3.85GB 10.82GB Fineweb-2, New CC
und_Samr 1.35M 158.99K 1.51M 64.06M 7.54M 71.59M 4.30GB 1.72GB 6.02GB Fineweb-2, New CC
ori_Orya 1.34M 145.91K 1.48M 128.69M 11.97M 140.66M 2.16GB 770.15MB 2.93GB MaLA
jav_Latn 1.26M 122.51K 1.38M 379.69M 35.26M 414.95M 1.96GB 587.75MB 2.55GB Fineweb-2, MaLA, New CC
yid_Hebr 1.25M 160.66K 1.41M 287.37M 36.14M 323.51M 2.84GB 1.30GB 4.14GB MaLA, New CC
und_Cans 1.23M 248.05K 1.48M 106.39M 21.43M 127.83M 3.48GB 2.77GB 6.25GB Fineweb-2, New CC
Table 8: Data Cleaning Statistics (part III): Comparison of document count, token count, disk size, and sources before and after data cleaning in DCAD-2000.
Lang Code Documents Tokens Disk Size Source
keep remove total keep remove total keep remove total
nya_Latn 1.21M 138.31K 1.34M 230.59M 26.29M 256.88M 1.34GB 437.81MB 1.78GB Fineweb-2, MaLA
hmn_Latn 1.20M 195.18K 1.40M 173.07M 28.59M 201.66M 1.08GB 543.90MB 1.63GB MaLA
tir_Ethi 1.20M 78.32K 1.28M 125.79M 8.16M 133.96M 1.15GB 290.56MB 1.44GB Fineweb-2, MaLA
uig_Arab 1.19M 78.60K 1.27M 513.72M 37.42M 551.15M 3.72GB 937.66MB 4.65GB Fineweb-2, MaLA, New CC
wln_Latn 1.18M 74.38K 1.25M 53.99M 3.61M 57.59M 520.40MB 78.21MB 598.61MB Fineweb-2, MaLA, New CC
und_Adlm 1.12M 194.29K 1.32M 43.63M 7.55M 51.18M 1.10GB 853.95MB 1.95GB Fineweb-2, New CC
und_Egyp 1.12M 190.50K 1.31M 97.41M 16.58M 113.99M 2.54GB 3.52GB 6.05GB Fineweb-2, New CC
und_Syrc 1.12M 115.88K 1.23M 42.71M 4.43M 47.14M 20.68GB 4.34GB 25.01GB Fineweb-2, New CC
swh_Latn 1.12M 82.67K 1.20M 449.92M 32.71M 482.63M 3.34GB 803.12MB 4.15GB Fineweb-2, MaLA
yor_Latn 1.12M 108.67K 1.22M 189.62M 18.77M 208.39M 1.08GB 304.95MB 1.38GB Fineweb-2, MaLA, New CC
uzn_Latn 1.03M 68.06K 1.10M 466.19M 30.78M 496.97M 4.03GB 1.03GB 5.06GB Fineweb-2, New CC
und_Mend 1.03M 293.72K 1.32M 16.58M 4.75M 21.33M 893.39MB 2.06GB 2.95GB Fineweb-2, New CC
xho_Latn 1.02M 88.44K 1.11M 168.59M 13.93M 182.52M 1.19GB 247.71MB 1.44GB Fineweb-2, MaLA
gla_Latn 1.01M 115.44K 1.13M 518.47M 76.34M 594.81M 2.03GB 904.76MB 2.94GB Fineweb-2, MaLA, New CC
bre_Latn 980.75K 86.36K 1.07M 134.68M 11.68M 146.37M 757.53MB 231.46MB 988.99MB Fineweb-2, MaLA, New CC
sot_Latn 917.37K 78.48K 995.85K 223.24M 17.82M 241.06M 1.09GB 283.15MB 1.37GB Fineweb-2, MaLA
nan_Latn 905.48K 86.68K 992.16K 26.58M 2.54M 29.12M 483.99MB 95.09MB 579.08MB Fineweb-2, MaLA
tel_Latn 898.42K 92.51K 990.93K 204.17M 21.27M 225.44M 843.51MB 444.92MB 1.29GB Fineweb-2, MaLA
bew_Latn 885.97K 99.33K 985.30K 370.27M 41.51M 411.78M 2.85GB 776.53MB 3.62GB Fineweb-2, New CC
smo_Latn 883.15K 83.25K 966.41K 241.45M 21.17M 262.62M 1.15GB 290.83MB 1.44GB Fineweb-2, MaLA
glk_Arab 876.52K 99.66K 976.18K 44.95M 5.30M 50.24M 630.38MB 171.44MB 801.82MB Fineweb-2, MaLA
che_Cyrl 875.25K 117.29K 992.54K 118.78M 15.18M 133.96M 1.05GB 346.83MB 1.40GB Fineweb-2, MaLA, New CC
orm_Latn 859.55K 77.40K 936.95K 35.46M 3.19M 38.65M 476.68MB 150.02MB 626.69MB MaLA
zho_Hani 840.53K 65.42K 905.95K 578.50M 46.68M 625.18M 2.67GB 980.93MB 3.65GB MaLA
haw_Latn 808.97K 88.12K 897.10K 227.68M 23.61M 251.29M 869.19MB 300.40MB 1.17GB Fineweb-2, MaLA
pnb_Arab 806.70K 71.03K 877.73K 133.55M 11.76M 145.31M 881.83MB 493.40MB 1.38GB Fineweb-2, MaLA, New CC
oci_Latn 760.65K 59.16K 819.82K 123.30M 10.54M 133.84M 706.68MB 193.69MB 900.37MB Fineweb-2, MaLA, New CC
und_Linb 735.07K 107.67K 842.75K 52.97M 7.76M 60.73M 6.30GB 997.90MB 7.30GB Fineweb-2, New CC
chv_Cyrl 731.68K 60.72K 792.40K 188.93M 16.35M 205.28M 1.10GB 361.84MB 1.46GB Fineweb-2, MaLA, New CC
kin_Latn 701.70K 67.29K 768.99K 197.65M 16.84M 214.49M 1.43GB 160.27MB 1.59GB Fineweb-2, MaLA
srp_Latn 630.88K 54.65K 685.53K 158.44M 13.19M 171.63M 775.01MB 209.48MB 984.49MB MaLA
und_Brai 590.10K 125.33K 715.43K 57.85M 12.29M 70.13M 1.94GB 1.30GB 3.24GB Fineweb-2, New CC
kaa_Cyrl 588.71K 48.01K 636.72K 1.08B 86.21M 1.16B 3.58GB 620.59MB 4.20GB Fineweb-2, MaLA
lug_Latn 570.88K 40.31K 611.19K 36.43M 2.65M 39.08M 344.92MB 85.21MB 430.13MB Fineweb-2, MaLA
und_Sgnw 567.29K 106.45K 673.74K 37.34M 7.01M 44.34M 1.40GB 1.11GB 2.50GB Fineweb-2, New CC
pcm_Latn 563.55K 80.45K 644.00K 135.97M 19.60M 155.57M 1.45GB 231.26MB 1.68GB Fineweb-2, MaLA
pbt_Arab 556.45K 36.70K 593.15K 273.04M 18.00M 291.04M 2.40GB 481.43MB 2.88GB Fineweb-2, MaLA
min_Latn 548.22K 32.98K 581.19K 28.26M 1.78M 30.04M 326.92MB 43.32MB 370.24MB Fineweb-2, MaLA
tuk_Latn 526.60K 48.40K 575.00K 211.69M 23.04M 234.74M 1.14GB 368.23MB 1.51GB Fineweb-2, MaLA
lim_Latn 526.45K 43.83K 570.28K 49.16M 4.85M 54.01M 338.07MB 70.26MB 408.33MB Fineweb-2, MaLA, New CC
und_Hung 520.10K 155.23K 675.33K 42.34M 12.64M 54.98M 1.94GB 2.32GB 4.25GB Fineweb-2, New CC
gsw_Latn 519.60K 64.76K 584.36K 171.13M 22.15M 193.28M 2.02GB 248.45MB 2.27GB Fineweb-2, MaLA, New CC
aze_Arab 481.85K 107.19K 589.05K 16.65M 3.70M 20.35M 283.94MB 125.40MB 409.33MB MaLA
kmr_Latn 473.75K 37.03K 510.79K 239.78M 19.24M 259.01M 1.64GB 366.13MB 2.01GB Fineweb-2, MaLA, New CC
roh_Latn 467.79K 40.88K 508.66K 59.96M 5.00M 64.96M 373.84MB 133.62MB 507.46MB Fineweb-2, MaLA, New CC
vec_Latn 451.53K 28.94K 480.47K 35.51M 2.41M 37.92M 248.96MB 70.25MB 319.21MB Fineweb-2, MaLA
san_Deva 426.60K 30.30K 456.90K 186.19M 14.19M 200.38M 1.37GB 884.42MB 2.25GB Fineweb-2, MaLA, New CC
und_Bali 422.49K 77.08K 499.57K 39.62M 7.23M 46.84M 1.19GB 662.91MB 1.85GB Fineweb-2, New CC
und_Nshu 419.71K 89.40K 509.11K 38.53M 8.21M 46.74M 993.06MB 1.28GB 2.27GB Fineweb-2, New CC
und_Modi 386.82K 67.33K 454.15K 52.58M 9.15M 61.73M 16.45GB 7.42GB 23.87GB Fineweb-2, New CC
gmh_Latn 383.58K 47.47K 431.05K 769.12M 95.18M 864.30M 5.51GB 1.42GB 6.93GB Fineweb-2, New CC
sco_Latn 382.19K 37.49K 419.69K 43.05M 4.46M 47.52M 357.63MB 98.46MB 456.10MB Fineweb-2, MaLA
nds_Latn 379.54K 44.24K 423.78K 79.45M 11.68M 91.13M 384.74MB 126.48MB 511.22MB Fineweb-2, MaLA, New CC
und_Lana 377.58K 110.80K 488.38K 47.55M 13.95M 61.50M 688.16MB 2.05GB 2.74GB Fineweb-2, New CC
azb_Arab 376.14K 24.16K 400.30K 81.10M 6.51M 87.61M 615.69MB 203.89MB 819.58MB Fineweb-2, MaLA, New CC
tsn_Latn 375.82K 23.43K 399.25K 24.79M 1.54M 26.33M 206.56MB 41.32MB 247.88MB Fineweb-2, MaLA
und_Mong 364.92K 51.36K 416.28K 78.04M 10.98M 89.03M 1.32GB 827.40MB 2.15GB Fineweb-2, New CC
sah_Cyrl 357.02K 24.17K 381.19K 110.13M 7.77M 117.89M 1.05GB 202.76MB 1.25GB MaLA, New CC
und_Ethi 351.77K 49.20K 400.97K 39.75M 5.56M 45.31M 1.23GB 1.08GB 2.31GB Fineweb-2, New CC
rus_Latn 349.61K 47.55K 397.17K 77.31M 10.54M 87.85M 755.00MB 485.49MB 1.24GB MaLA
pri_Latn 348.99K 27.20K 376.20K 142.27M 11.09M 153.36M 2.15GB 505.82MB 2.66GB Fineweb-2, New CC
und_Hebr 345.20K 46.87K 392.07K 17.42M 2.36M 19.78M 548.23MB 461.10MB 1.01GB Fineweb-2, New CC
mon_Latn 344.80K 46.68K 391.48K 31.56M 4.27M 35.84M 180.12MB 271.24MB 451.37MB MaLA
pap_Latn 339.80K 22.62K 362.42K 127.89M 8.52M 136.41M 678.73MB 223.10MB 901.83MB Fineweb-2, MaLA
tgk_Latn 337.95K 48.39K 386.35K 26.44M 3.79M 30.22M 198.08MB 219.19MB 417.27MB MaLA
plt_Latn 330.57K 28.23K 358.80K 118.46M 8.02M 126.48M 951.31MB 189.98MB 1.14GB Fineweb-2, MaLA
lmo_Latn 324.18K 29.25K 353.43K 41.37M 4.09M 45.46M 230.80MB 58.92MB 289.72MB Fineweb-2, MaLA, New CC
bod_Tibt 318.52K 34.22K 352.75K 252.06M 28.33M 280.39M 3.37GB 998.77MB 4.37GB MaLA, New CC
und_Saur 315.78K 73.82K 389.60K 15.26M 3.57M 18.83M 398.55MB 489.07MB 887.62MB Fineweb-2, New CC
yue_Hani 300.49K 34.04K 334.53K 9.02M 1.03M 10.06M 790.86MB 161.03MB 951.90MB Fineweb-2, MaLA, New CC
Table 9: Data Cleaning Statistics (part IV): Comparison of document count, token count, disk size, and sources before and after data cleaning in DCAD-2000.
Lang Code Documents Tokens Disk Size Source
keep remove total keep remove total keep remove total
bar_Latn 270.31K 30.79K 301.10K 92.48M 12.46M 104.94M 318.42MB 142.36MB 460.78MB Fineweb-2, MaLA
und_Thaa 262.00K 37.56K 299.56K 7.68M 1.10M 8.78M 391.60MB 263.21MB 654.81MB Fineweb-2, New CC
und_Dupl 258.90K 53.06K 311.96K 14.14M 2.90M 17.04M 752.58MB 502.95MB 1.26GB Fineweb-2, New CC
arg_Latn 258.20K 22.52K 280.72K 29.97M 3.05M 33.02M 207.88MB 43.58MB 251.45MB Fineweb-2, MaLA, New CC
pms_Latn 258.13K 20.25K 278.38K 23.55M 1.86M 25.41M 172.17MB 39.05MB 211.22MB Fineweb-2, MaLA, New CC
hif_Latn 254.95K 37.47K 292.41K 220.19M 38.74M 258.93M 779.02MB 879.71MB 1.66GB Fineweb-2, MaLA
und_Thai 254.35K 47.64K 301.99K 47.88M 8.97M 56.85M 868.70MB 325.83MB 1.19GB Fineweb-2, New CC
und_Runr 252.18K 39.00K 291.18K 154.68M 23.92M 178.61M 1.25GB 3.28GB 4.52GB Fineweb-2, New CC
und_Vaii 243.47K 93.27K 336.73K 71.28M 27.31M 98.59M 513.30MB 1.88GB 2.39GB Fineweb-2, New CC
vol_Latn 241.22K 23.73K 264.95K 12.26M 1.28M 13.54M 126.45MB 27.79MB 154.24MB Fineweb-2, MaLA, New CC
und_Glag 237.68K 72.07K 309.75K 20.38M 6.18M 26.56M 476.61MB 951.96MB 1.43GB Fineweb-2, New CC
nrm_Latn 234.99K 31.99K 266.97K 71.12M 9.68M 80.80M 654.26MB 233.97MB 888.23MB Fineweb-2, MaLA
aeb_Arab 230.69K 32.19K 262.88K 51.79M 7.23M 59.01M 641.42MB 232.91MB 874.33MB Fineweb-2, New CC
kat_Latn 229.64K 46.98K 276.62K 37.42M 7.66M 45.08M 247.34MB 365.42MB 612.76MB MaLA
ido_Latn 222.87K 22.62K 245.49K 15.65M 1.48M 17.13M 131.86MB 35.81MB 167.67MB Fineweb-2, MaLA, New CC
kal_Latn 220.32K 17.35K 237.67K 76.08M 6.03M 82.11M 371.13MB 202.28MB 573.42MB Fineweb-2, MaLA
pam_Latn 219.65K 22.53K 242.18K 21.42M 2.45M 23.87M 129.69MB 37.16MB 166.84MB Fineweb-2, MaLA
und_Khmr 216.99K 36.25K 253.24K 10.97M 1.83M 12.80M 473.35MB 417.98MB 891.34MB Fineweb-2, New CC
lus_Latn 206.91K 16.42K 223.33K 66.59M 5.16M 71.75M 387.16MB 114.21MB 501.37MB Fineweb-2, MaLA
und_Mymr 204.74K 27.30K 232.03K 5.63M 751.18K 6.39M 283.14MB 249.17MB 532.31MB Fineweb-2, New CC
und_Tibt 201.49K 32.84K 234.33K 15.44M 2.52M 17.95M 970.52MB 505.25MB 1.48GB Fineweb-2, New CC
und_Dsrt 198.00K 37.90K 235.90K 4.47M 855.49K 5.32M 248.83MB 562.92MB 811.75MB Fineweb-2, New CC
und_Geor 196.35K 49.50K 245.85K 22.22M 5.60M 27.83M 374.55MB 629.81MB 1.00GB Fineweb-2, New CC
new_Deva 187.27K 16.23K 203.49K 23.86M 2.07M 25.93M 302.85MB 89.56MB 392.41MB Fineweb-2, MaLA, New CC
und_Mroo 186.14K 22.85K 208.99K 6.42M 788.69K 7.21M 2.43GB 335.38MB 2.77GB Fineweb-2, New CC
sme_Latn 184.43K 14.88K 199.30K 42.27M 3.53M 45.80M 318.80MB 92.35MB 411.15MB Fineweb-2, MaLA
und_Bopo 181.71K 24.45K 206.16K 30.63M 4.12M 34.75M 3.45GB 890.68MB 4.35GB Fineweb-2, New CC
nso_Latn 175.98K 9.66K 185.64K 18.89M 1.08M 19.97M 111.81MB 32.30MB 144.10MB Fineweb-2, MaLA
und_Armn 168.06K 46.69K 214.75K 33.05M 9.18M 42.24M 347.17MB 515.75MB 862.92MB Fineweb-2, New CC
und_Mtei 166.92K 19.64K 186.57K 48.49M 5.71M 54.20M 795.85MB 570.91MB 1.37GB Fineweb-2, New CC
scn_Latn 162.55K 10.71K 173.26K 18.07M 1.48M 19.55M 125.29MB 25.45MB 150.75MB Fineweb-2, MaLA, New CC
ina_Latn 159.80K 16.99K 176.79K 13.62M 1.49M 15.11M 104.31MB 28.02MB 132.33MB Fineweb-2, MaLA, New CC
lld_Latn 154.22K 24.98K 179.20K 8.00M 1.25M 9.25M 90.50MB 16.81MB 107.31MB Fineweb-2, MaLA
und_Khar 153.37K 40.04K 193.41K 6.75M 1.76M 8.52M 250.30MB 182.38MB 432.67MB Fineweb-2, New CC
hyw_Armn 142.71K 12.89K 155.60K 60.77M 5.52M 66.29M 863.69MB 174.18MB 1.04GB Fineweb-2, MaLA
und_Deva 141.13K 26.07K 167.20K 35.96M 6.64M 42.60M 255.99MB 1.53GB 1.79GB Fineweb-2, New CC
abk_Cyrl 139.72K 12.86K 152.58K 7.73M 671.84K 8.40M 100.31MB 14.57MB 114.88MB Fineweb-2, MaLA
und_Brah 138.03K 22.72K 160.75K 7.85M 1.29M 9.15M 273.71MB 243.75MB 517.47MB Fineweb-2, New CC
bpy_Beng 135.66K 9.50K 145.16K 9.30M 766.18K 10.07M 141.90MB 28.27MB 170.17MB Fineweb-2, MaLA, New CC
bew_Cyrl 133.83K 13.49K 147.32K 3.37M 339.68K 3.71M 74.12MB 15.81MB 89.93MB MaLA
lin_Latn 133.64K 8.68K 142.32K 16.04M 1.37M 17.41M 115.63MB 32.27MB 147.89MB Fineweb-2, MaLA
und_Bhks 131.90K 27.03K 158.93K 3.93M 805.58K 4.74M 190.96MB 154.63MB 345.59MB Fineweb-2, New CC
oss_Cyrl 128.06K 13.97K 142.03K 84.80M 9.56M 94.36M 390.43MB 167.02MB 557.45MB Fineweb-2, MaLA, New CC
tgk_Arab 127.77K 14.97K 142.75K 11.61M 1.36M 12.97M 104.11MB 55.51MB 159.62MB MaLA
szl_Latn 127.60K 10.33K 137.93K 8.52M 738.21K 9.25M 89.99MB 12.81MB 102.80MB Fineweb-2, MaLA
mww_Latn 122.30K 10.22K 132.52K 98.37M 8.22M 106.59M 536.48MB 104.20MB 640.68MB Fineweb-2, New CC
sdh_Arab 120.04K 14.20K 134.24K 35.26M 4.50M 39.76M 466.52MB 136.99MB 603.52MB Fineweb-2, MaLA
und_Hmnp 118.87K 12.33K 131.20K 6.83M 708.37K 7.54M 436.28MB 151.81MB 588.09MB Fineweb-2, New CC
srd_Latn 118.78K 8.14K 126.92K 15.38M 1.23M 16.61M 119.77MB 24.18MB 143.95MB Fineweb-2, MaLA
mhr_Cyrl 118.77K 12.58K 131.35K 30.71M 3.17M 33.88M 278.82MB 75.27MB 354.09MB Fineweb-2, MaLA, New CC
ydd_Hebr 117.78K 7.28K 125.06K 73.71M 4.55M 78.26M 879.66MB 120.71MB 1.00GB Fineweb-2, MaLA
diq_Latn 117.09K 11.78K 128.87K 9.75M 962.88K 10.71M 75.44MB 16.34MB 91.79MB Fineweb-2, MaLA, New CC
und_Telu 115.91K 30.83K 146.74K 9.00M 2.39M 11.40M 409.30MB 426.99MB 836.29MB Fineweb-2, New CC
que_Latn 114.28K 23.93K 138.21K 4.28M 896.83K 5.18M 57.76MB 33.84MB 91.59MB MaLA, New CC
run_Latn 114.03K 9.29K 123.32K 24.63M 1.97M 26.60M 218.56MB 39.33MB 257.89MB Fineweb-2, MaLA
hsb_Latn 112.76K 9.95K 122.71K 25.10M 2.04M 27.14M 153.09MB 23.81MB 176.90MB Fineweb-2, MaLA, New CC
wol_Latn 108.94K 11.08K 120.02K 11.76M 1.37M 13.13M 95.99MB 29.75MB 125.74MB Fineweb-2, MaLA
rmy_Latn 108.23K 21.11K 129.34K 284.56M 55.71M 340.26M 2.54GB 98.95MB 2.64GB Fineweb-2, MaLA
und_Phag 107.75K 17.58K 125.34K 3.41M 556.36K 3.97M 141.68MB 93.31MB 234.99MB Fineweb-2, New CC
und_Merc 107.52K 38.04K 145.56K 7.61M 2.69M 10.30M 215.43MB 472.23MB 687.66MB Fineweb-2, New CC
urd_Latn 106.75K 12.60K 119.35K 139.19M 16.43M 155.63M 312.70MB 140.28MB 452.98MB Fineweb-2, New CC
kiu_Latn 106.48K 10.36K 116.84K 36.53M 3.76M 40.29M 289.67MB 193.73MB 483.39MB Fineweb-2, MaLA
cak_Latn 106.28K 6.64K 112.92K 6.08M 438.75K 6.52M 66.17MB 10.86MB 77.03MB Fineweb-2, MaLA
ilo_Latn 106.18K 7.83K 114.01K 28.61M 2.06M 30.67M 143.69MB 37.61MB 181.30MB Fineweb-2, MaLA, New CC
und_Kali 105.87K 24.33K 130.19K 1.39M 318.99K 1.71M 105.22MB 91.44MB 196.66MB Fineweb-2, New CC
und_Plrd 104.31K 21.07K 125.38K 5.47M 1.10M 6.57M 214.53MB 225.25MB 439.77MB Fineweb-2, New CC
und_Orya 104.03K 26.52K 130.56K 10.14M 2.59M 12.73M 299.43MB 387.64MB 687.07MB Fineweb-2, New CC
und_Lisu 101.47K 20.05K 121.52K 24.00M 4.74M 28.74M 204.23MB 527.19MB 731.42MB Fineweb-2, New CC
und_Hmng 101.02K 23.34K 124.36K 5.37M 1.24M 6.61M 153.20MB 196.99MB 350.19MB Fineweb-2, New CC